### v1

# Coherence-Assisted Entanglement Activation During Pure Dephasing Evolution

Abstract
We investigate the system-environment entanglement activation during pure dephasing evolution, also called controlled unitary. A lower bound of generated entanglement is formulated with the initial coherence in the system and using relative entropy measure. For a qubit-environment case, this lower becomes a closed form by replacing conditional entropy with the overlap on the states in environment. Applying this result to the spin-boson model, we illustrate how the lower bound evolves as a function of time multiplied cutoff frequency, shown plots. The different behaviors of lower bounds with respect to Ohmicity, s=2 and s=3, corresponding to Markovian and non-Markovian evolution. In addition, we show that the decoherence in the system throughout the pure dephasing evolution represents another lower bound and upper bound of generated entanglement, indicating that larger entanglement generation leads more decoherence.
1 Introduction
Entanglement Horodecki et al. (2009) has been the subject of quite amount of researches in last decades due to its significance not only for quantum information and computation, but for the foundations of quantum mechanics as well. Understandably, the vast majority of entanglement studies focus on the purposeful entanglement generation, manipulation, and preservation. However, there are processes where entanglement can be generated spontaneously. Open systems dynamics (see e.g. Schlosshauer (2007); Breuer and Petruccione (2002) for a modern introduction) is one of such examples, where the interaction between the system and its environment leads in general to entanglement creation.
Until recently Roszak and Cywiński (2015); Roszak (2018, 2020); Rzepkowski and Roszak (2020) this system-environment entanglement has not been systematically studied using modern methods of quantum information theory, being mostly treated as a mere source of decoherence. However, with the emergence and the gradual acceptance of a new paradigm in open systems, given by the quantum Darwinism idea Zurek (2009) and its refinements such as Spectrum Broadcast Structures and strong Darwinism (see Korbicz (2021) for a review), the environment has been recognized as a carrier of valuable information about the system. Therefore, studying system-environment entanglement is important not only for understanding decoherence processes, since entanglement in general increases decoherence efficiency, see e.g. Takahashi et al. (2022) for recent take on this question, but also to know how environment acquires system-related information.
One can illustrate system-environment entanglement generation on a paradigmatic example of the, so called, controlled unitary evolution, alternatively called also a measurement-type or a pure dephasing evolution, and defined by:
$|s_{i}\rangle|E_{0}\rangle\rightarrow|s_{i}\rangle|E_{i}\rangle,$ (1)
in some preferred basis of the system $|s_{i}\rangle$, called the pointer basis  Zurek (1981, 1982, 2003). This basis is specified by the nature of the interaction, e.g. the dipole approximation for the electromagnetic interaction, leading to the spin-boson model Schlosshauer (2007); Breuer and Petruccione (2002). Whenever there is a superposition present in the initial state $|\psi_{S}\rangle=\sum_{i}a_{i}|s_{i}\rangle$, the resulting state $|\psi^{\prime}_{SE}\rangle$ is entangled:
$|\psi_{S}\rangle|E_{0}\rangle\rightarrow|\psi^{\prime}_{SE}\rangle=\sum_{i}a_{i}|s_{i}\rangle|E_{i}\rangle$ (2)
while the reduced subsystem $\textrm{Tr}_{E}[|\psi^{\prime}_{SE}\rangle\langle\psi^{\prime}_{SE}|]$ lost the initial coherence in $|\psi_{S}\rangle$. In the modern language of the resource theories Chitambar and Gour (2019), this an example of the resource interconversion, where one resource, coherence in the pointer basis, gets interconverted into another - entanglement  Piani et al. (2011); Chuan et al. (2012); Streltsov et al. (2015); Killoran et al. (2016); Qiao et al. (2018); Regula et al. (2018). Let us remark here that an abstract coherence theory finds its natural application in open systems with pure dephasing evolution as the latter contains a natural, preferred basis.
However easy the pure state interconversion (2) looks like, the situation gets much more complicated when considering mixed states, especially if one wants to quantify the amount of the generated entanglement. Assuming that the initial state of the system is pure, Roszak et al.  Roszak and Cywiński (2015); Roszak (2018) recently found necessary and sufficient separability conditions for the system-environment split and have developed own operational measure of the system-environment entanglement Roszak (2020); Rzepkowski and Roszak (2020). However, it is still unknown how to generalize their methods to mixed initial states of the system.
Previous studies of the conversion from coherence to entanglement showed that the entanglement generated by incoherent operations between coherent state in the system and incoherent pure state in environment is upper bounded by the initial coherence in the system respect to relative entropy measure Streltsov et al. (2015). However, for given Hamiltonian, the corresponding pure dephasing evolution is not incoherent operation generally and we also take the environmental state as arbitrary, their finding does not hold in our case. In pure dephasing evolution, therefore, it is not clear how entanglement generation is quantified through the initial coherence or decoherence in the system.
In this work, we study entanglement generation for mixed states of the system. We derive two lower bounds and one upper bound for the system-environment entanglement as measured by the relative entropy of entanglement $E_{r}$ Vedral et al. (1997); Vedral and Plenio (1998). Following the simple intuition (2) and its developments Piani et al. (2011); Chuan et al. (2012); Streltsov et al. (2015); Killoran et al. (2016); Qiao et al. (2018); Regula et al. (2018), the first lower bound is given in terms of the initial coherence of the system with respect to the natural pointer basis defined by the pure dephasing evolution, diminished by the remaining uncertainty about the system after the best measurement on the environment.
The second lower bound and the upper bound are quantified with the decoherence in the system, following the intuition that entanglement generation requires not only the initial coherence but the decoherence the system after the evolution which is also suggested by (2).
Regarding the first lower bound, for two-dimensional central systems, i.e. for qubits, we derive a stronger result in given by a closed formula that links entanglement with the distinguishability of the environment states. In particular, this allows to study entanglement dynamics for open qubits, understood as the real-time monitoring of the minimum amount of entanglement produced during the evolution. As an illustration, we apply our results to the spin-boson model (see e.g. Schlosshauer (2007) and the references therein). We study the dependence of the minimum generated entanglement on various model parameters e.g. the temperature and Ohmicity. Varying the latter through the well-known Markovian/non-Markovian transition, we see the traces of non-Markovianity in the dynamics of entanglement; cf. Rivas et al. (2010).
2 Preliminaries
Pure dephasing evolution is defined by the following Hamiltonian:
$H=H_{S}+H_{E}+\sum_{i=0}^{N-1}|i\rangle\langle i|\otimes V_{i}.$ (3)
$H_{S}$ is the self Hamiltonian of the system which here we assume to be commuting with the interaction, $H_{S}=\sum_{i}\varepsilon_{i}|i\rangle\langle i|$. The case of a more general $H_{S}$ is out of the scope of this work as we concentrate on the pure dephasing models. $H_{E}$ and $V_{i}$’s are arbitrary environmental observables. (3) is the most general Hamiltonian that generates evolution of the type (2). The unitary evolution $U$ is given by:
$U(t)=\sum_{i}|i\rangle\langle i|\otimes U_{E}^{i}(t),\quad U_{E}^{i}(t)=e^{-(\varepsilon_{i}+H_{E}+V_{i})t},$ (4)
which is the controlled-unitary or measurement-type operation in the quantum information terminology. The basis $|i\rangle$ is the pointer basis. Since observables $V_{i}$ are arbitrary, in general non-commuting, there does not exist a common diagonalizing basis, i.e. an incoherent basis for the environment, which makes the whole unitary operator incoherent. We will review the definition of incoherent operations below.
We will assume that the system and the environment are initially product, i.e. $\rho_{SE}(0)=\rho_{S}\otimes\rho_{E}$. This assumption is justified here by the fact that we are interested in the build-up of correlations, thus we want to initiate the system in a completely uncorrelated state, so that any produced correlations are due to the interaction. Applying the unitary operator defined in (4), we obtain the following state $\sigma_{SE}$ after the interaction:
$\sigma_{SE}=U(\rho_{S}\otimes\rho_{E})U^{\dagger}=\sum_{i,j}|i\rangle\langle i|\rho_{S}|j\rangle\langle j|\otimes U^{i}_{E}\rho_{E}{U^{j}_{E}}^{\dagger}$ (5)
In order to evaluate the amount of $SE$ entanglement, we will use the relative entropy of entanglement $E_{r}$ Vedral et al. (1997); Vedral and Plenio (1998), defined as:
$E_{r}(\rho_{SE})=\min_{\xi_{SE}\in Sep}S(\rho_{SE}||\xi_{SE})$ (6)
where $S(\rho||\xi)=\textrm{Tr}[\rho(\log\rho-\log\xi)]$ is the quantum relative entropy and the minimization is taken over all separable states $\xi_{SE}=\sum_{i}p_{i}\xi_{S}^{i}\otimes\xi_{E}^{i}$. This is a proper measure of entanglement Plenio and Virmani (2007), with a clear operational interpretation as the exponential bound on asymptotic probability of error of confusing $\rho_{AB}$ with the nearest separable state via POVM measurements, see e.g. Vedral (2002). The minimization in definition is the reason why it is difficult to evaluate the value directly. We employ the following lemma to bypass this problem:
Lemma 1.
Plenio et al. (2000) The relative entropy of entanglement is bounded from below by the negative conditional entropy, i.e. for any bipartite state $\rho_{AB}$ we have:
$E_{r}(\rho_{AB})\geq\max\{S(\rho_{A})-S(\rho_{AB}),S(\rho_{B})-S(\rho_{AB})\}$ (7)
where $S(\rho)$ is the von Neumann entropy: $S(\rho)=-\textrm{Tr}[\rho\log\rho]$.
We will also need elements of the quantum coherence theory; see e.g.  Baumgratz et al. (2014). Since quantum coherence is a basis dependent resource, we need to specify which basis is incoherent. In principle, one can develop a theory with any incoherent basis. However, since we study a concrete family of open dynamics, given by (3), the pointer basis of the central system $|i\rangle$ is the natural candidate here because if the system loses coherence due to the interaction, decoheres to this basis. We do not specify the incoherent basis for environment because the coherence in the environment does not matter in our analysis as we will show below. In the language of the coherence theory, an incoherent state is simply a diagonal state in the incoherent basis, which is of course nothing else but a decohered state in the open systems/quantum foundations terminology. We denote by $\mathcal{I}$ the set of all incoherent states. In a resource theory, of which coherence theory is an example, an operation which does not create resource is called a free operation respect to that resource. Various free operations for coherence are proposed for different physical motivations Chitambar and Gour (2016). The most standard one is incoherent operation whose Kraus operators $\{K_{n}\}$ satisfy $\frac{K_{n}\rho K_{n}^{\dagger}}{\textrm{Tr}[K_{n}\rho K_{n}^{\dagger}]}\in\mathcal{I}$ for all $n$ and any incoherent state $\rho\in\mathcal{I}$. One can now define the relative entropy of coherence $C_{r}$ in analogy to the relative entropy of entanglement (6) Baumgratz et al. (2014):
$C_{r}(\rho)=\min_{\sigma\in\mathcal{I}}S(\rho||\sigma)=S(\Delta(\rho))-S(\rho)$ (8)
where $\Delta(\rho)$ is called the dephasing channel, defined as $\Delta(\rho)=\sum_{i}|i\rangle\langle i|\rho|i\rangle\langle i|$.
Another ingredient that we will need is a measure of state distinguishability. Out of the possible measures Fuchs and van de Graaf (1999), we will use the state fidelity $B(\rho,\sigma)$, also called generalized overlap, defined as:
$B(\rho,\sigma)=$ $\min_{\textrm{POVM}:\{M_{i}\}}\sum_{i}\sqrt{\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]}=\textrm{Tr}\sqrt{\rho^{\frac{1}{2}}\sigma\rho^{\frac{1}{2}}}$ (9)
where $\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]$ is the probability that one obtains the same value $i$ from the measurement on $\rho$ and $\sigma$. Vanishing state fidelity is equivalent to a perfect distinguishability of the two states by POVM’s. This justifies the use of the state fidelity as a measure of distinguishability. This measure does not have as clear operational meaning as e.g. the quantum Chernoff bound Audenaert et al. (2007), related similarly to the quantum relative entropy to the asymptotic discrimination error of two states, but it is much easier to work with due to the compact form (9) and it also provides both an upper and a lower bound for the quantum Chernoff information.
Discriminating states from an ensemble =$\mathcal{E}=\{p_{i},\rho_{i}\}$ is a more difficult task and its efficiency can be quantified by the, so called, accessible information $I_{acc}(\mathcal{E})$. It is defined as the maximum of the Shannon (classical) mutual information $I(i:m)$ Cover and Thomas (2006) between the ensemble index $i$ and the measurement result $m$, taken over all POVM’s $\{M_{m}\}$ on $\rho=\sum_{i}p_{i}\rho_{i}$. The formal expression is given by:
$I_{acc}=\max_{POVM:\{M_{m}\}}I(i:m),$ (10)
where $I(i:m)$ is defined using the conditional probabilities of measuring $m$ while the sate was $\rho_{i}$: $p(m|i)=\textrm{Tr}(M_{m}\rho_{i})$.
For two-element ensembles, the two notions of discrimination are connected by the following Lemma, which will be central for our reasoning:
Lemma 2.
Rahul and Ashwin (2007)[Lemma 3.6] Let =$\mathcal{E}=\{p,\rho_{0},(1-p),\rho_{1}\}$ be a two-element ensemble of quantum states of arbitrary finite dimensions. Then:
$I_{acc}(\mathcal{E})\geq h(p)-2\sqrt{p(1-p)}B(\rho_{0},\rho_{1})$ (11)
where $h$ is the binary Shannon entropy: $h(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$.
3 Bounds on the system-environment entanglement generation
In the pure dephasing, open evolution, described in the Introduction, the system’s information is transferred to the environment through the controlled-unitary interaction (4). This interaction can be interpreted as the environment implementing a measurement in the pointer basis $\{|i\rangle\}$ and storing the information in its state. Indeed, as one can easily prove, the state of the environment in the course of the evolution becomes:
$\sigma_{E}(t)=\textrm{Tr}_{S}[U(t)\rho_{S}\otimes\rho_{E}U(t)^{\dagger}]=\sum_{i}p_{i}\rho_{i}(t),$ (12)
where $p_{i}=\langle i|\rho_{S}|i\rangle$ are the initial pointer probabilities and the conditional states of the environment $\rho_{i}(t)=U_{E}^{i}(t)\rho_{E}{U_{E}^{i}}(t)^{\dagger}$ encode, in a generally imperfect way, the information about the state of the central system. One can now try to recover this information by implementing a POVM measurement $\{M_{m}\}$ on the environment in order to extract the value of $i$, encoded by the environment in the course of the evolution. One may view this as an indirect measurement of the system through the environment, common in the macroscopic world; cf. the studies of objectivity in quantum mechanics in e.g. Zurek (2009); Korbicz (2021). The uncertainty of this indirect measurement is represented by Shannnon conditional entropy $H(i|m)$ Cover and Thomas (2006). From above observations, it is natural to think the less this uncertainty is, the stronger the system and the environment are correlated. On the other hand, the more is there initial coherence in the pointer basis $\{|i\rangle\}$, the more entanglement should be generated Killoran et al. (2016). Indeed this is so as the following Lemma shows:
Lemma 3.
For any POVM measurement $\{M_{m}\}$ on the environment, the relative entropy of entanglement, generated by the pure dephasing evolution $U$, is bounded from below by:
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq C_{r}(\rho_{S})-H(i|m).$ (13)
Proof.
Applying Lemma 1 to $\sigma_{SE}=U\rho_{S}\otimes\rho_{E}U^{\dagger}$, we have
[…]
$S(\sigma_{E})-S(\sigma_{SE})\}.$ (14)
We focus on finding a lower bound for $S(\sigma_{E})-S(\sigma_{SE})$ because $S(\sigma_{E})$ can be transformed to the conditional entropy or the overlap of environmental states through lemma 2, leading us to a nice result represented in corollary 1. We will also consider a lower bound for $S(\sigma_{S})-S(\sigma_{SE})$ later. We first observe that:
$S(\sigma_{E})-S(\sigma_{SE})=S(\sigma_{E})-S(\rho_{S})-S(\rho_{E})$ (15)
where we used fact that the unitary evolution preserves the von Neumann entropy and it is additive for product states. Notice (15) can be negative while the relative entanglement should be always non-negative.
Next, let us recall the Holevo bound on the $I_{acc}$ associated with (12):
$I_{acc}\leq S(\sigma_{E})-\sum_{i}\langle i|\rho_{S}|i\rangle S(U_{i}\rho_{E}U_{i}^{\dagger}),$ (16)
from which it follows that:
$S(\sigma_{E})\geq I_{acc}(\sigma_{E})+S(\rho_{E}),$ (17)
since $S(U_{i}\rho_{E}U_{i}^{\dagger})=S(\rho_{E})$. From the definition of the Shannon mutual information $I(x:y)=H(x)-H(x|y)$, where $H(x)$ and $H(x|y)$ are the classical Shannon and relative entropies respectively, the accessible information $I_{acc}$ can be represented as follows.
$I_{acc}=\max_{POVM:\{M_{m}\}}I(i:m)=\max_{POVM:\{M_{m}\}}\left[H(i)-H(i|m)\right].$ (18)
Index $i$ is a random variable, distributed with the initial pointer probabilities $p_{i}$, as follows from (12), or equivalently with the eigenvalues of the fully decohered state $\Delta(\rho_{S})=\sum_{i}|i\rangle\langle i|\rho_{S}|i\rangle\langle i|$. Therefore we may write $H(i)=S(\Delta(\rho_{S}))$. Since there is a maximization over all POVM in the definition (18), for an arbitrary POVM $\{M_{m}\}$ we will have:
$I_{acc}\geq H(i)-H(i|m)=S(\Delta(\rho_{S}))-H(i|m)$ (19)
Now we recover the statement by substituting (17) and (19) into (15):
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq$ $S(\sigma_{E})-S(\rho_{S})-S(\rho_{E})$
$\geq$ $I_{acc}-S(\rho_{S})$ (20)
$\geq$ $S(\Delta(\rho_{S}))-H(i|m)-S(\rho_{S})$
$=$ $C_{r}(\rho_{S})-H(i|m),$
where in the final line we used the fact (8). ∎
The above Lemma is valid for any finete dimensionality of the central system and the environment. The price to pay for this generality is the presence of the arbitrary POVM in the bound (15). This arbitrariness can be removed when the central system is a qubit. Then, there ensemble (12) consists of only two states $\rho_{0}(t)=U_{0}(t)\rho_{E}U_{0}(t)^{\dagger}$, $\rho_{1}(t)=U_{1}(t)\rho_{E}U_{1}(t)^{\dagger}$ with probabilities $p=\langle 0|\rho_{S}|0\rangle$ and $1-p=\langle 1|\rho_{S}|1\rangle$. We can then apply the Lemma 2 and obtain:
Corollary 1.
If the central system is two-dimensional, then:
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})$
$\geq$ $C_{r}(\rho_{S})-2\sqrt{p(1-p)}B(\rho_{0}(t),\rho_{1}(t)).$ (21)
The proof follows immediately from the application of Lemma 2 in (20).
Let us look closely at Lemma 3 and give some interpretations. The second term $H(i|m)$ on the right hand side is the uncertainty about the system’s state measured by the environment after our best measurement on environment. Non zero $H(i|m)$ means that we cannot conclude the system’s value in the pointer basis even if we could use any POVM on the environment. Reversely, if $H(i|m)$ becomes zero, the information about the system is transferred fully to environment and entanglement must be there unless system’s coherence is zero, which means the system is fully classical state. After all, when the system has more coherence in pointer bases, more entanglement is required to know the value of system certainly. One can also re-arrange the inequality (13) and obtain the lower bound on the uncertainty about the system’s state in terms of initial coherences and the generated entanglement:
$\max_{POVM:\{M_{m}\}}H(i|m)\geq C_{r}(\rho_{S})-E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger}).$ (22)
Similar interpretation can be applied to the qubit version (1). If $p(1-p)=0$ one knows the system’s state with certainty. Notice that $C_{r}(\rho_{S})=0$ in this case. Also if $B(\rho_{0}(t),\rho_{1}(t))=0$, it implies that $\rho_{0}(t)$, $\rho_{1}(t)$ are perfectly distinguishable, which means maximum information has been transferred to the environment.
Note that the bounds (13),(1) are not tight and can even become negative as we have mentioned in the proof and will see in an example in the next section. In that case of course the bounds trivialize as $0\leq E_{r}$ always. Nevertheless, despite the lack of optimality, they provide a valuable information about the open evolution, certifying that at least a given amount of entanglement has been generated. This is particularly interesting when the right hand sides of (13),(1) are large, which implies a large amount of created entanglement.
Finally, we will give an upper bound and another lower bound which we mentioned in the proof of lemma 3 on the generated entanglement. From a previous research Streltsov et al. (2015), an upper bound on entanglement generated by incoherent operations with initially coherent system $\rho_{S}$ and an incoherent pure ancilla $|0\rangle\langle 0|_{E}$ is given by the initial coherence in system, for any measure of entanglement $E_{D}$ and coherence $C_{D}$ corresponding to a contractive distance $D$:
$E_{D}(\Lambda[\rho_{S}\otimes|0\rangle\langle 0|_{E}])\leq C_{D}(\rho_{S})$ (23)
where $\Lambda$ is any incoherent operation on the bipartite system. When we specify $D$ as the relative entropy, which we have been using throughout this work, the above upper bound becomes tight.
However, since the corresponding unitary evolution $U$ for given Hamiltonian of (3) is not incoherent in general regardless of incoherent basis for environment and we assume the environmental state arbitrary, the above upper bound is not valid in our scenario. Also, considering our intuitive picture of system-environment interaction where coherence is converted into entanglement, it is interesting to connect the generated entanglement not only to the initial coherence $C_{r}(\rho_{S})$ as we did in lemma 3 for lower bound, but to the decoherence, the difference of coherence between the initial and final states in system as we demonstrate below.
Lemma 4.
Let $D_{S}$ be the decoherence in the system defined as $D_{S}=C_{r}(\rho_{S})-C_{r}(\sigma_{S})$. Then the relative entropy of entanglement generated by pure dephasing evolution $U$ is bounded by below and above as:
$D_{S}-S(\rho_{E})\leq E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\leq D_{S}+H(i).$ (24)
Proof.
Let us prove the left inequality first. From lemma 1, $E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq S(\sigma_{S})-S(\sigma_{SE})$. The invariancy of entropy over unitary evolution and the fact $S(\Delta(\sigma_{S}))=S(\Delta(\rho_{S}))$ lead us to the followings.
$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq S(\sigma_{S})-S(\sigma_{SE})$
$=$ $S(\sigma_{S})-S(\rho_{S})-S(\rho_{E})$
$=$ $S(\sigma_{S})-S(\Delta(\rho_{S}))+S(\Delta(\rho_{S}))-S(\rho_{S})-S(\rho_{E})$
$=$ $C_{r}(\rho_{S})-C_{r}(\sigma_{S})-S(\rho_{E})$ (25)
Moving to the right hand side, from the definition of the relative entropy of entanglement: $E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\leq S(\sigma_{SE}||\sigma_{S}\otimes\sigma_{E})=I(S:E)=S(\sigma_{S})+S(\sigma_{E})-S(\sigma_{SE})=S(\sigma_{S})+S(\sigma_{E})-S(\rho_{S})-S(\rho_{E})$, since by definition $\sigma_{SE}=U\rho_{S}\otimes\rho_{E}U^{\dagger}$ so that $S(\sigma_{SE})=S(\rho_{S}\otimes\rho_{E})=S(\rho_{S})+S(\rho_{E})$. We also observe that $S(\sigma_{E})=S(\sum_{i}p_{i}\rho_{i})\leq S(\rho_{E})+H(i)$, which follows from the well known property of the von Neumann entropy: $S(\sum_{i}p_{i}\rho_{i})\leq\sum p_{i}S(\rho_{i})+H(i)$. Substituting this into the above inequality, we prove the statement. ∎
Lemma 4 fits the intuition that larger system-environment entanglement leads to more decoherence quantified by $C_{r}(\rho_{S})-C_{r}(\sigma_{S})$. In other words, decoherence in the system $D_{S}$ is lower bounded by generated entanglement.
4 Case study: The spin-boson model
In this Section, we illustrate our results on the spin-boson model, which is one of the most popular open systems models; see e.g. Schlosshauer (2007); Breuer and Petruccione (2007), and the references therein for an introduction. In its most used form, it describes a two-level central system, a qubit, interacting with a collection of bosonic modes (or harmonic oscillators), constituting the environment; for an example of higher dimensional central systems see e.g. Kiciński and Korbicz (2021). The dynamics of the model is given by the following Hamiltonian:
$H$ $=\sum_{\boldsymbol{k}}\omega_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}a_{\boldsymbol{k}}+\sigma_{z}\otimes\sum_{\boldsymbol{k}}\left(g_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}+g^{*}_{\boldsymbol{k}}a_{\boldsymbol{k}}\right),$ (26)
where for the purpose of this example, we have suppressed the self-Hamiltonian of the central system, or assumed it to be commuting with the interaction, as otherwise the evolution would not be of the pure dephasing form (4). Here, $a^{\dagger}_{\boldsymbol{k}}$, $a_{\boldsymbol{k}}$ are the annihilation and creation operators of the mode $\boldsymbol{k}$ of the environment, $\sigma_{z}$ is the $z$ axis Pauli matrix acting in the space of the central system and $g_{\boldsymbol{k}}$ are the coupling constants. The pointer basis here is the eigenbasis of $\sigma_{z}$, which follows immediately from the interaction term in (26).
We assume the environment to be thermal at temperature $T$, i.e. $\rho_{E}=e^{-\beta H_{E}}/Z$, where we choose the units so that $\beta=1/T$, and the central system to be initially in the following mixed state:
$\rho_{S}=(1-\alpha)|+\rangle\langle+|+\alpha\frac{\mathbbm{I}}{2}.$ (27)
It is a mixture of a state that is maximally coherent w.r.t. the pointer basis and the maximally mixed state $\mathbbm{I}/2$, and $\alpha$ is the mixedness parameter, $0\leq\alpha\leq 1$. Such a choice of the initial state will allow us to control the initial coherence in the pointer basis.
Since the central system is two-dimensional, we will use the bound (1). The initial coherence $C_{r}(\rho_{S})$ can be easily calculated and is given by the binary entropy:
$C_{r}(\rho_{S})=1-h\left(\frac{\alpha}{2}\right)$ (28)
Interestingly, the fidelity $B(\rho_{0}(t),\rho_{1}(t))$ can also be calculated analytically for the popular Ohmic spectral density:
$J(\omega)=\frac{\omega^{s}}{\Lambda^{s-1}}\bm{e}^{-\omega/\Lambda},$ (29)
where $\Lambda$ is the cutoff frequency and $s$ is the Ohmicity parameter. The value $s=3$ is a known threshold, distinguishing Markovian evolution for $s<3$ from non-Markovian for $s\geq 3$; see e.g. Addis et al. (2014). In Lampo et al. (2017) the analytical expression for $B$ was obtained for the case of the Ohmicity parameter $s>1$. It consists of two parts, vacuum, i.e. temperature-independent, and thermal:
$\ln B(t)=$ $\ln B_{vac}(t)+\ln B_{th}(t),$ (30)
$\frac{1}{2}\ln B_{vac}(t)=$ $-\wp(s-1)\left[1-\frac{\cos[(s-1)\arctan(\Lambda t)]}{(1+\Lambda^{2}t^{2})^{\frac{s-1}{2}}}\right],$ (31)
$\frac{1}{4}\ln B_{th}(t)=$ $\left(-\frac{T}{2\Lambda}\right)^{s-1}\bigg{\{}\Psi^{(s-2)}\left(1+\frac{T}{2\Lambda}\right)-\Psi^{(s-2)}\left(\frac{1}{2}+\frac{T}{2\Lambda}\right)$
[…]
where c.c. means complex conjugate of the last two terms in the parenthesis, $\wp(s)$ is the Euler gamma function, and $\Psi^{m}(z)$ is the polygamma function defined as:
$\Psi^{m}(z)=\sum_{k=0}^{\infty}\frac{(-1)^{m+1}m!}{(z+k)^{m+1}}.$ (33)
Figure 1: Lower bound (1) for the spin-boson model as a function of time for a different mixedness parameter $\alpha$ (cf. (27)) and different Ohmicity: a) $s=2$ corresponding to Markovian evolution; b) $s=3$ corresponding to non-Markovian evolution. The ratio of the thermal to the cutoff energy is set to $T/\Lambda=1$ and the time is measured in the cutoff units. The insert shows the peak and thus the non-monotonic behavior of the plots (with different vertical scale for each plot). On all plots the parts, where the bound becomes negative were replaced by the trivial bound equal to zero.
[…]
In Fig. 1, we keep fixed the ratio of the thermal to the cutoff energy, $T/\Lambda=1$, and plot (1) for different mixedness $\alpha$ and Ohmicity $s$ parameters. As expected from the (1), less mixed, i.e. more coherent, states with lower $\alpha$ lead to much more entanglement generation. More interestingly, there is a slight quantitative change of behavior between $s=2$ and $s=3$, corresponding to the transition between the Markovian and non-Markovian evolution respectively. While for $s=2$ the behavior is monotonic in time, for $s=3$ there is a broad peak around $\Lambda t=2$, regardless the value of $\alpha$ (depicted on the insert). This peak is more visible in Fig. 2, where we fix the mixedness $\alpha=1/2$ and vary the energy ratio $T/\Lambda$. The non-monotonic behavior of the entanglement bound for $s=3$ bears certain similarity to non-Markovianity measures, where a non-monotonic behavior of some marker function certifies information back-flow and hence the non-Markovianity of the evolution. Especially close seems the measure from Hall et al. (2014), where the non-Markovianity is detected via non-monotonicity of the decoherence rate, which in turn derives from system-environment correlations.
We also notice from Fig. 1 that the higher the $T/\Lambda$ ratio, the higher the entanglement bound, which means that if the cutoff frequency $\Lambda$ is fixed, higher temperatures guarantee stronger entanglement between the system and the environment. This in turn leads to a stronger decoherence and explains the well known fact why the system need to be kept cold in order to maintain the coherence.
5 Conclusions
We studied generation of entanglement between the system of interest and its environment for a broad class of so called pure dephasing evolution. We extended the previous results  Roszak and Cywiński (2015); Roszak (2018), obtained for pure initial states of the system, to the case of mixed states. We combined methods and results from coherence theory, state discrimination, and open systems evolution to obtain a lower bound on the generated entanglement, measured by the relative entropy.
Our second result describes another lower bound and upper bound for generated entanglement in terms of the difference of coherence in the system before and after the evolution. These bounds also can be seen as an extension of the result obtained in Streltsov et al. (2015). Although both the lower bound and upper bound contains the extra terms other than decoherence, we successfully prove that larger entanglement indicates more decoherence in pure dephasing evolution with arbitrary environmental state and vice versa.
Regarding the first result, its lower bound character is a price to pay for the generalization to mixed states. However, it certifies that at least the given amount of entanglement was created during the evolution. The first result is the strongest for the two-dimensional central system, where we obtained a compact formula. For arbitrary dimension, our bound still contains a certain optimization. As an example, we calculated analytically and studied numerically our bound for the popular spin-boson system, rediscovering the well-known traces of the non-Markovianity but in the entanglement dynamics. Obviously, more systems can be studied, especially that in some of them the fidelities has already been calculated Korbicz et al. (2014); Tuziemski et al. (2019); Kwiatkowski et al. (2021); Kiciński and Korbicz (2021)
Acknowledgements
We would like to thank K. Roszak for getting us interested in the topic and for fruitful discussions and comments. We acknowledge the financial support of the Polish National Science Foundation (NCN) through the grants Opus 2019/35/B/ST2/01896 and Quant-Era “Quantum Coherence Activation By Open Systems and Environments” QuCABOoSE 2023/05/Y/ST2/00139.

### v2

@@
-Coherence-Assisted Entanglement Activation During Pure Dephasing Evolution
+Coherence-Assisted Entanglement Activation During Open Evolution
 Abstract
-We investigate the system-environment entanglement activation during pure dephasing evolution, also called controlled unitary.
-A lower bound of generated entanglement is formulated with the initial coherence in the system and using relative entropy measure.
-For a qubit-environment case, this lower becomes a closed form by replacing conditional entropy with the overlap on the states in environment.
-Applying this result to the spin-boson model, we illustrate how the lower bound evolves as a function of time multiplied cutoff frequency, shown plots.
-The different behaviors of lower bounds with respect to Ohmicity, s=2 and s=3, corresponding to Markovian and non-Markovian evolution.
-In addition, we show that the decoherence in the system throughout the pure dephasing evolution represents another lower bound and upper bound of generated entanglement, indicating that larger entanglement generation leads more decoherence.
+The fact that in the course of an open evolution there is, in a vast majority of cases, a creation of system-environment entanglement out of the initial coherences, has been known for many years.
+However, how much entanglement can be created under general assumptions has not been analyzed yet.
+In this work, we investigate system-environment entanglement generation under a broad and important class of interactions known as pure dephasing interactions.
+Using completely general initials states, we derive a series of bounds on the amount of the created entanglement, connecting it to the initial system coherences as well as to the decohrence strength, making quantitatively precise the statement that larger entanglement leads to more decoherence.
+Interestingly, we obtain lower bounds, certifying minimal amount of entanglement that can be produced.
+As an illustration, we apply our results to the spin-boson model, where a qubit interacts with a bosonic bath, showing time evolution of the entanglement and the Markovian/non-Markovian crossover.
+Our results help to better understand and quantify the complicated correlations, produced during an open evolution, which will hopefully lead to better/new decoherence mitigation techniques.
 1 Introduction
-Entanglement Horodecki et al.
-(2009) has been the subject of quite amount of researches in last decades due to its significance not only for quantum information and computation, but for the foundations of quantum mechanics as well.
+Entanglement [1] has been the subject of quite amount of researches in last decades due to its significance not only for quantum information and computation, but for the foundations of quantum mechanics as well.
 Understandably, the vast majority of entanglement studies focus on the purposeful entanglement generation, manipulation, and preservation.
-However, there are processes where entanglement can be generated spontaneously.
+However, there are processes where entanglement can be generated or activated spontaneously.
 Open systems dynamics (see e.g.
-Schlosshauer (2007); Breuer and Petruccione (2002) for a modern introduction) is one of such examples, where the interaction between the system and its environment leads in general to entanglement creation.
-Until recently Roszak and Cywiński (2015); Roszak (2018, 2020); Rzepkowski and Roszak (2020) this system-environment entanglement has not been systematically studied using modern methods of quantum information theory, being mostly treated as a mere source of decoherence.
-However, with the emergence and the gradual acceptance of a new paradigm in open systems, given by the quantum Darwinism idea Zurek (2009) and its refinements such as Spectrum Broadcast Structures and strong Darwinism (see Korbicz (2021) for a review), the environment has been recognized as a carrier of valuable information about the system.
+[2, 3] for a modern introduction) is one of such examples, where the interaction between the system and its environment leads in general to entanglement creation out of initial system coherences.
+Until recently [4, 5, 6, 7] this system-environment entanglement has not been systematically studied using modern methods of quantum information theory, being mostly treated as a mere source of decoherence.
+However, with the emergence and the gradual acceptance of a new paradigm in open systems, given by the quantum Darwinism idea [8] and its refinements such as Spectrum Broadcast Structures and strong Darwinism (see e.g.
+[9] for a review), the environment has been recognized as a carrier of valuable information about the system.
 Therefore, studying system-environment entanglement is important not only for understanding decoherence processes, since entanglement in general increases decoherence efficiency, see e.g.
-Takahashi et al.
-(2022) for recent take on this question, but also to know how environment acquires system-related information.
-One can illustrate system-environment entanglement generation on a paradigmatic example of the, so called, controlled unitary evolution, alternatively called also a measurement-type or a pure dephasing evolution, and defined by:
+[10] for recent take on this question, but also to know how environment acquires system-related information.
+One can illustrate the system-environment entanglement activation on a paradigmatic example of the, so called, pure dephasing evolution, also known as controlled-unitary evolution, alternatively called also a measurement-type evolution:
 $|s_{i}\rangle|E_{0}\rangle\rightarrow|s_{i}\rangle|E_{i}\rangle,$ (1)
-in some preferred basis of the system $|s_{i}\rangle$, called the pointer basis  Zurek (1981, 1982, 2003).
+in some preferred basis of the system $|s_{i}\rangle$, called the pointer basis  [11, 12, 13].
 This basis is specified by the nature of the interaction, e.g.
-the dipole approximation for the electromagnetic interaction, leading to the spin-boson model Schlosshauer (2007); Breuer and Petruccione (2002).
-Whenever there is a superposition present in the initial state $|\psi_{S}\rangle=\sum_{i}a_{i}|s_{i}\rangle$, the resulting state $|\psi^{\prime}_{SE}\rangle$ is entangled:
+the dipole approximation for the electromagnetic interaction, leading to the spin-boson model [2, 3].
+Whenever there is a coherence present in the initial state, $|\psi_{S}\rangle=\sum_{i}a_{i}|s_{i}\rangle$, the resulting state $|\psi^{\prime}_{SE}\rangle$ is obviously entangled:
 $|\psi_{S}\rangle|E_{0}\rangle\rightarrow|\psi^{\prime}_{SE}\rangle=\sum_{i}a_{i}|s_{i}\rangle|E_{i}\rangle$ (2)
 while the reduced subsystem $\textrm{Tr}_{E}[|\psi^{\prime}_{SE}\rangle\langle\psi^{\prime}_{SE}|]$ lost the initial coherence in $|\psi_{S}\rangle$.
-In the modern language of the resource theories Chitambar and Gour (2019), this an example of the resource interconversion, where one resource, coherence in the pointer basis, gets interconverted into another - entanglement  Piani et al.
-(2011); Chuan et al.
-(2012); Streltsov et al.
-(2015); Killoran et al.
-(2016); Qiao et al.
-(2018); Regula et al.
-(2018).
+In the modern language of the resource theories [14], this an example of the resource interconversion, where one resource, coherence in the pointer basis, gets interconverted into another - entanglement  [15, 16, 17, 18, 19, 20].
 Let us remark here that an abstract coherence theory finds its natural application in open systems with pure dephasing evolution as the latter contains a natural, preferred basis.
@@
 Assuming that the initial state of the system is pure, Roszak et al.
-Roszak and Cywiński (2015); Roszak (2018) recently found necessary and sufficient separability conditions for the system-environment split and have developed own operational measure of the system-environment entanglement Roszak (2020); Rzepkowski and Roszak (2020).
+[4, 5] recently found necessary and sufficient separability conditions for the system-environment split and have developed own operational measure of the system-environment entanglement [6, 7].
 However, it is still unknown how to generalize their methods to mixed initial states of the system.
-Previous studies of the conversion from coherence to entanglement showed that the entanglement generated by incoherent operations between coherent state in the system and incoherent pure state in environment is upper bounded by the initial coherence in the system respect to relative entropy measure Streltsov et al.
-(2015).
+Previous studies of the conversion from coherence to entanglement showed that the entanglement generated by incoherent operations between coherent state in the system and incoherent pure state in environment is upper bounded by the initial coherence in the system respect to relative entropy measure [17].
 However, for given Hamiltonian, the corresponding pure dephasing evolution is not incoherent operation generally and we also take the environmental state as arbitrary, their finding does not hold in our case.
@@
 In this work, we study entanglement generation for mixed states of the system.
-We derive two lower bounds and one upper bound for the system-environment entanglement as measured by the relative entropy of entanglement $E_{r}$ Vedral et al.
-(1997); Vedral and Plenio (1998).
-Following the simple intuition (2) and its developments Piani et al.
-(2011); Chuan et al.
-(2012); Streltsov et al.
-(2015); Killoran et al.
-(2016); Qiao et al.
-(2018); Regula et al.
-(2018), the first lower bound is given in terms of the initial coherence of the system with respect to the natural pointer basis defined by the pure dephasing evolution, diminished by the remaining uncertainty about the system after the best measurement on the environment.
+We derive two lower bounds and one upper bound for the system-environment entanglement as measured by the relative entropy of entanglement $E_{r}$ [21, 22].
+Following the simple intuition (2) and its developments [15, 16, 17, 18, 19, 20], the first lower bound is given in terms of the initial coherence of the system with respect to the natural pointer basis defined by the pure dephasing evolution, diminished by the remaining uncertainty about the system after the best measurement on the environment.
 The second lower bound and the upper bound are quantified with the decoherence in the system, following the intuition that entanglement generation requires not only the initial coherence but the decoherence the system after the evolution which is also suggested by (2).
@@
 As an illustration, we apply our results to the spin-boson model (see e.g.
-Schlosshauer (2007) and the references therein).
+[2] and the references therein).
 We study the dependence of the minimum generated entanglement on various model parameters e.g.
@@
 Varying the latter through the well-known Markovian/non-Markovian transition, we see the traces of non-Markovianity in the dynamics of entanglement; cf.
-Rivas et al.
-(2010).
+[23].
 2 Preliminaries
@@
 Applying the unitary operator defined in (4), we obtain the following state $\sigma_{SE}$ after the interaction:
-$\sigma_{SE}=U(\rho_{S}\otimes\rho_{E})U^{\dagger}=\sum_{i,j}|i\rangle\langle i|\rho_{S}|j\rangle\langle j|\otimes U^{i}_{E}\rho_{E}{U^{j}_{E}}^{\dagger}$ (5)
-In order to evaluate the amount of $SE$ entanglement, we will use the relative entropy of entanglement $E_{r}$ Vedral et al. (1997); Vedral and Plenio (1998), defined as:
+$\sigma_{SE}=U(\rho_{S}\otimes\rho_{E})U^{\dagger}$
+$=\sum_{i,j}|i\rangle\langle i|\rho_{S}|j\rangle\langle j|\otimes U^{i}_{E}\rho_{E}{U^{j}_{E}}^{\dagger}$ (5)
+In order to evaluate the amount of $SE$ entanglement, we will use the relative entropy of entanglement $E_{r}$ [21, 22], defined as:
 $E_{r}(\rho_{SE})=\min_{\xi_{SE}\in Sep}S(\rho_{SE}||\xi_{SE})$ (6)
 where $S(\rho||\xi)=\textrm{Tr}[\rho(\log\rho-\log\xi)]$ is the quantum relative entropy and the minimization is taken over all separable states $\xi_{SE}=\sum_{i}p_{i}\xi_{S}^{i}\otimes\xi_{E}^{i}$.
-This is a proper measure of entanglement Plenio and Virmani (2007), with a clear operational interpretation as the exponential bound on asymptotic probability of error of confusing $\rho_{AB}$ with the nearest separable state via POVM measurements, see e.g.
-Vedral (2002).
+This is a proper measure of entanglement [24], with a clear operational interpretation as the exponential bound on asymptotic probability of error of confusing $\rho_{AB}$ with the nearest separable state via POVM measurements, see e.g.
+[25].
 The minimization in definition is the reason why it is difficult to evaluate the value directly.
@@
 Lemma 1.
-Plenio et al. (2000) The relative entropy of entanglement is bounded from below by the negative conditional entropy, i.e. for any bipartite state $\rho_{AB}$ we have:
+[26] The relative entropy of entanglement is bounded from below by the negative conditional entropy, i.e. for any bipartite state $\rho_{AB}$ we have:
 $E_{r}(\rho_{AB})\geq\max\{S(\rho_{A})-S(\rho_{AB}),S(\rho_{B})-S(\rho_{AB})\}$ (7)
@@
 We will also need elements of the quantum coherence theory; see e.g.
-Baumgratz et al.
-(2014).
+[27].
 Since quantum coherence is a basis dependent resource, we need to specify which basis is incoherent.
@@
 In a resource theory, of which coherence theory is an example, an operation which does not create resource is called a free operation respect to that resource.
-Various free operations for coherence are proposed for different physical motivations Chitambar and Gour (2016).
+Various free operations for coherence are proposed for different physical motivations [28].
 The most standard one is incoherent operation whose Kraus operators $\{K_{n}\}$ satisfy $\frac{K_{n}\rho K_{n}^{\dagger}}{\textrm{Tr}[K_{n}\rho K_{n}^{\dagger}]}\in\mathcal{I}$ for all $n$ and any incoherent state $\rho\in\mathcal{I}$.
-One can now define the relative entropy of coherence $C_{r}$ in analogy to the relative entropy of entanglement (6) Baumgratz et al.
-(2014):
+One can now define the relative entropy of coherence $C_{r}$ in analogy to the relative entropy of entanglement (6) [27]:
 $C_{r}(\rho)=\min_{\sigma\in\mathcal{I}}S(\rho||\sigma)=S(\Delta(\rho))-S(\rho)$ (8)
@@
 Another ingredient that we will need is a measure of state distinguishability.
-Out of the possible measures Fuchs and van de Graaf (1999), we will use the state fidelity $B(\rho,\sigma)$, also called generalized overlap, defined as:
-$B(\rho,\sigma)=$ $\min_{\textrm{POVM}:\{M_{i}\}}\sum_{i}\sqrt{\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]}=\textrm{Tr}\sqrt{\rho^{\frac{1}{2}}\sigma\rho^{\frac{1}{2}}}$ (9)
+Out of the possible measures [29], we will use the state fidelity $B(\rho,\sigma)$, also called generalized overlap, defined as:
+$B(\rho,\sigma)=$ $\min_{\textrm{POVM}:\{M_{i}\}}\sum_{i}\sqrt{\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]}$
+$=$ $\textrm{Tr}\sqrt{\rho^{\frac{1}{2}}\sigma\rho^{\frac{1}{2}}}$ (9)
 where $\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]$ is the probability that one obtains the same value $i$ from the measurement on $\rho$ and $\sigma$.
@@
 This measure does not have as clear operational meaning as e.g.
-the quantum Chernoff bound Audenaert et al.
-(2007), related similarly to the quantum relative entropy to the asymptotic discrimination error of two states, but it is much easier to work with due to the compact form (9) and it also provides both an upper and a lower bound for the quantum Chernoff information.
+the quantum Chernoff bound [30], related similarly to the quantum relative entropy to the asymptotic discrimination error of two states, but it is much easier to work with due to the compact form (2) and it also provides both an upper and a lower bound for the quantum Chernoff information.
 Discriminating states from an ensemble =$\mathcal{E}=\{p_{i},\rho_{i}\}$ is a more difficult task and its efficiency can be quantified by the, so called, accessible information $I_{acc}(\mathcal{E})$.
-It is defined as the maximum of the Shannon (classical) mutual information $I(i:m)$ Cover and Thomas (2006) between the ensemble index $i$ and the measurement result $m$, taken over all POVM’s $\{M_{m}\}$ on $\rho=\sum_{i}p_{i}\rho_{i}$.
+It is defined as the maximum of the Shannon (classical) mutual information $I(i:m)$ [31] between the ensemble index $i$ and the measurement result $m$, taken over all POVM’s $\{M_{m}\}$ on $\rho=\sum_{i}p_{i}\rho_{i}$.
 The formal expression is given by:
@@
 Lemma 2.
-Rahul and Ashwin (2007)[Lemma 3.6] Let =$\mathcal{E}=\{p,\rho_{0},(1-p),\rho_{1}\}$ be a two-element ensemble of quantum states of arbitrary finite dimensions. Then:
+[32][Lemma 3.6] Let =$\mathcal{E}=\{p,\rho_{0},(1-p),\rho_{1}\}$ be a two-element ensemble of quantum states of arbitrary finite dimensions. Then:
 $I_{acc}(\mathcal{E})\geq h(p)-2\sqrt{p(1-p)}B(\rho_{0},\rho_{1})$ (11)
@@
 the studies of objectivity in quantum mechanics in e.g.
-Zurek (2009); Korbicz (2021).
-The uncertainty of this indirect measurement is represented by Shannnon conditional entropy $H(i|m)$ Cover and Thomas (2006).
+[8, 9].
+The uncertainty of this indirect measurement is represented by Shannnon conditional entropy $H(i|m)$ [31].
 From above observations, it is natural to think the less this uncertainty is, the stronger the system and the environment are correlated.
-On the other hand, the more is there initial coherence in the pointer basis $\{|i\rangle\}$, the more entanglement should be generated Killoran et al.
-(2016).
+On the other hand, the more is there initial coherence in the pointer basis $\{|i\rangle\}$, the more entanglement should be generated [18].
 Indeed this is so as the following Lemma shows:
@@
 From the definition of the Shannon mutual information $I(x:y)=H(x)-H(x|y)$, where $H(x)$ and $H(x|y)$ are the classical Shannon and relative entropies respectively, the accessible information $I_{acc}$ can be represented as follows.
-$I_{acc}=\max_{POVM:\{M_{m}\}}I(i:m)=\max_{POVM:\{M_{m}\}}\left[H(i)-H(i|m)\right].$ (18)
+$I_{acc}=$ $\max_{POVM:\{M_{m}\}}I(i:m)$
+$=$ $\max_{POVM:\{M_{m}\}}\left[H(i)-H(i|m)\right].$ (18)
 Index $i$ is a random variable, distributed with the initial pointer probabilities $p_{i}$, as follows from (12), or equivalently with the eigenvalues of the fully decohered state $\Delta(\rho_{S})=\sum_{i}|i\rangle\langle i|\rho_{S}|i\rangle\langle i|$.
@@
 Finally, we will give an upper bound and another lower bound which we mentioned in the proof of lemma 3 on the generated entanglement.
-From a previous research Streltsov et al.
-(2015), an upper bound on entanglement generated by incoherent operations with initially coherent system $\rho_{S}$ and an incoherent pure ancilla $|0\rangle\langle 0|_{E}$ is given by the initial coherence in system, for any measure of entanglement $E_{D}$ and coherence $C_{D}$ corresponding to a contractive distance $D$:
+From a previous research [17], an upper bound on entanglement generated by incoherent operations with initially coherent system $\rho_{S}$ and an incoherent pure ancilla $|0\rangle\langle 0|_{E}$ is given by the initial coherence in system, for any measure of entanglement $E_{D}$ and coherence $C_{D}$ corresponding to a contractive distance $D$:
 $E_{D}(\Lambda[\rho_{S}\otimes|0\rangle\langle 0|_{E}])\leq C_{D}(\rho_{S})$ (23)
@@
 ∎
-Lemma 4 fits the intuition that larger system-environment entanglement leads to more decoherence quantified by $C_{r}(\rho_{S})-C_{r}(\sigma_{S})$.
-In other words, decoherence in the system $D_{S}$ is lower bounded by generated entanglement.
+Lemma 4 fits the intuition that larger system-environment entanglement leads to more decoherence as quantified by $D_{S}$.
+More precisely, we show that $D_{S}$ is lower bounded by the generated entanglement:
+$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})-H(i)\leq D_{S}.$ (26)
 4 Case study: The spin-boson model
 In this Section, we illustrate our results on the spin-boson model, which is one of the most popular open systems models; see e.g.
-Schlosshauer (2007); Breuer and Petruccione (2007), and the references therein for an introduction.
+[2, 3], and the references therein for an introduction.
 In its most used form, it describes a two-level central system, a qubit, interacting with a collection of bosonic modes (or harmonic oscillators), constituting the environment; for an example of higher dimensional central systems see e.g.
-Kiciński and Korbicz (2021).
+[33].
 The dynamics of the model is given by the following Hamiltonian:
-$H$ $=\sum_{\boldsymbol{k}}\omega_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}a_{\boldsymbol{k}}+\sigma_{z}\otimes\sum_{\boldsymbol{k}}\left(g_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}+g^{*}_{\boldsymbol{k}}a_{\boldsymbol{k}}\right),$ (26)
+$H$ $=\sum_{\boldsymbol{k}}\omega_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}a_{\boldsymbol{k}}+\sigma_{z}\otimes\sum_{\boldsymbol{k}}\left(g_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}+g^{*}_{\boldsymbol{k}}a_{\boldsymbol{k}}\right)$ (27)
 where for the purpose of this example, we have suppressed the self-Hamiltonian of the central system, or assumed it to be commuting with the interaction, as otherwise the evolution would not be of the pure dephasing form (4).
@@
 $J(\omega)=\frac{\omega^{s}}{\Lambda^{s-1}}\bm{e}^{-\omega/\Lambda},$ (30)
-where $\Lambda$ is the cutoff frequency and $s$ is the Ohmicity parameter.
+where $\Lambda$ is the cut-off frequency and $s$ is the Ohmicity parameter.
 The value $s=3$ is a known threshold, distinguishing Markovian evolution for $s<3$ from non-Markovian for $s\geq 3$; see e.g.
-Addis et al.
-(2014).
-In Lampo et al.
-(2017) the analytical expression for $B$ was obtained for the case of the Ohmicity parameter $s>1$.
+[34].
+In [35] the analytical expression for $B$ was obtained for the case of the Ohmicity parameter $s>1$.
 It consists of two parts, vacuum, i.e.
@@
 The non-monotonic behavior of the entanglement bound for $s=3$ bears certain similarity to non-Markovianity measures, where a non-monotonic behavior of some marker function certifies information back-flow and hence the non-Markovianity of the evolution.
-Especially close seems the measure from Hall et al.
-(2014), where the non-Markovianity is detected via non-monotonicity of the decoherence rate, which in turn derives from system-environment correlations.
+Especially close seems the measure from [36], where the non-Markovianity is detected via non-monotonicity of the decoherence rate, which in turn derives from system-environment correlations.
 We also notice from Fig.
@@
 5 Conclusions
-We studied generation of entanglement between the system of interest and its environment for a broad class of so called pure dephasing evolution.
-We extended the previous results  Roszak and Cywiński (2015); Roszak (2018), obtained for pure initial states of the system, to the case of mixed states.
-We combined methods and results from coherence theory, state discrimination, and open systems evolution to obtain a lower bound on the generated entanglement, measured by the relative entropy.
-Our second result describes another lower bound and upper bound for generated entanglement in terms of the difference of coherence in the system before and after the evolution.
-These bounds also can be seen as an extension of the result obtained in Streltsov et al.
-(2015).
-Although both the lower bound and upper bound contains the extra terms other than decoherence, we successfully prove that larger entanglement indicates more decoherence in pure dephasing evolution with arbitrary environmental state and vice versa.
-Regarding the first result, its lower bound character is a price to pay for the generalization to mixed states.
-However, it certifies that at least the given amount of entanglement was created during the evolution.
-The first result is the strongest for the two-dimensional central system, where we obtained a compact formula.
-For arbitrary dimension, our bound still contains a certain optimization.
-As an example, we calculated analytically and studied numerically our bound for the popular spin-boson system, rediscovering the well-known traces of the non-Markovianity but in the entanglement dynamics.
-Obviously, more systems can be studied, especially that in some of them the fidelities has already been calculated Korbicz et al.
-(2014); Tuziemski et al.
-(2019); Kwiatkowski et al.
-(2021); Kiciński and Korbicz (2021)
+We studied generation of entanglement between the system of interest and its environment for a broad class of so called pure dephasing evolutions, also known as measurement-type interactions.
+In this part of the work, we extended previous results  [4, 5] obtained for pure initial states of the system, to the case of mixed states.
+We combined methods and results from coherence theory, state discrimination, and open systems evolution to obtain a series of bounds on the amount of the generated entanglement as measured by the relative entropy.
+Firstly, and most importantly, we obtain lower bounds, which certify that at least the given amount of entanglement was created during the evolution out of a given amount of the initial coherence.
+The lower bound character of this result is a price to pay for the generalization to arbitrary, mixed states of the system.
+For qubit as the central system, where we obtained a closed expression, connecting amount of entanglement to the state distinguishability of the environmental states.
+This makes an interesting link to certain studies of quantum Darwinism [9].
+For an arbitrary dimension, our bound still contains a certain optimization.
+As an example, we calculated analytically and studied numerically the lower bound for the popular spin-boson system, rediscovering the well-known traces of the non-Markovianity but in the entanglement dynamics.
+Obviously, more systems can be analyzed, especially that for some of them the relevant state distinguishability has already been quantified [37, 38, 39, 33].
+Next, we derived both lower and upper bounds for the entanglement in terms of the difference of coherence in the system before and after the evolution, which is a measure of the decoherence efficiency.
+This proves in an elegant and general way that stronger entanglement leads to more decoherence and vice versa, more decoherence is accompanied with more entanglement generation.
+These bounds also can be seen as an extension of the result obtained in [17].
+Summarizing, our results allow to better characterize and understand the correlations that pure dephasing interaction can produce under general initial conditions.
+Because of the ubiquitous character of the pure dephasing interaction in open systems [2, 3], our results are of a fairly universal character across the field of open systems and decoherence theory.
+Moreover, since quantum measurement theory is largely based on the same type of interaction, we characterize how much entanglement is generated during a measurement process, thus linking to foundations of quantum mechanics.
+On the other hand, a possible practical importance of our studies may lie in the development of non-standard decoherence-mitigation techniques, which rely on the system-environment correlations (see e.g.
+[40] for an example of such a technique).
 Acknowledgements

### v5

@@
 Abstract
-The fact that in the course of an open evolution there is, in a vast majority of cases, a creation of system-environment entanglement out of the initial coherences, has been known for many years.
+The fact that in the course of an open evolution there is, in a vast majority of cases, the creation of system-environment entanglement out of the initial coherence, has been known for many years.
 However, how much entanglement can be created under general assumptions has not been analyzed yet.
-In this work, we investigate system-environment entanglement generation under a broad and important class of interactions known as pure dephasing interactions.
-Using completely general initials states, we derive a series of bounds on the amount of the created entanglement, connecting it to the initial system coherences as well as to the decohrence strength, making quantitatively precise the statement that larger entanglement leads to more decoherence.
-Interestingly, we obtain lower bounds, certifying minimal amount of entanglement that can be produced.
+In this work, we investigate system-environment entanglement generation under a broad and important class of interactions known as pure dephasing interactions and for arbitrary initial conditions.
+Our main results are the following bounds on the relative entropy of entanglement $E_{r}(\sigma_{SE})$ after evolution: $C_{r}(\rho_{S})-H(\mathcal{I}|\mathcal{M})\leq E_{r}(\sigma_{SE})\leq C_{r}(\rho_{S})-C_{r}(\sigma_{S})+H(\mathcal{I})$, connecting it on one side to the initial relative entropy of coherence, $C_{r}(\rho_{S})$, and information extractable from the environment, $H(\mathcal{I}|\mathcal{M})$, and on the other side to the decoherence strength, $C_{r}(\rho_{S})-C_{r}(\sigma_{S})$, making quantitatively precise the statement that larger entanglement leads to more decoherence.
 As an illustration, we apply our results to the spin-boson model, where a qubit interacts with a bosonic bath, showing time evolution of the entanglement and the Markovian/non-Markovian crossover.
-Our results help to better understand and quantify the complicated correlations, produced during an open evolution, which will hopefully lead to better/new decoherence mitigation techniques.
-1 Introduction
-Entanglement [1] has been the subject of quite amount of researches in last decades due to its significance not only for quantum information and computation, but for the foundations of quantum mechanics as well.
+Our results help better understand and quantify the complicated correlations, produced during an open evolution, which will hopefully lead to better and new decoherence mitigation techniques.
+I Introduction
+Entanglement [1] has been the subject of a significant amount of research in the last decades due to its significance not only for quantum information and computation, but for the foundations of quantum mechanics as well.
 Understandably, the vast majority of entanglement studies focus on the purposeful entanglement generation, manipulation, and preservation.
 However, there are processes where entanglement can be generated or activated spontaneously.
-Open systems dynamics (see e.g.
-[2, 3] for a modern introduction) is one of such examples, where the interaction between the system and its environment leads in general to entanglement creation out of initial system coherences.
-Until recently [4, 5, 6, 7] this system-environment entanglement has not been systematically studied using modern methods of quantum information theory, being mostly treated as a mere source of decoherence.
+Open system dynamics (see e.g.
+[2, 3] for a modern introduction) is one such example, where the interaction between the system and its environment leads in general to entanglement creation out of initial system coherences.
+Until recently [4, 5, 6, 7], this system-environment entanglement had not been systematically studied using modern methods of quantum information theory, having been mostly treated as a mere source of decoherence.
 However, with the emergence and the gradual acceptance of a new paradigm in open systems, given by the quantum Darwinism idea [8] and its refinements such as Spectrum Broadcast Structures and strong Darwinism (see e.g.
@@
 [10] for recent take on this question, but also to know how environment acquires system-related information.
-One can illustrate the system-environment entanglement activation on a paradigmatic example of the, so called, pure dephasing evolution, also known as controlled-unitary evolution, alternatively called also a measurement-type evolution:
+One can illustrate the system-environment entanglement activation on a paradigmatic example of the so-called pure dephasing evolution, also known as controlled-unitary evolution, alternatively called also a measurement-type evolution:
 $|s_{i}\rangle|E_{0}\rangle\rightarrow|s_{i}\rangle|E_{i}\rangle,$ (1)
@@
 while the reduced subsystem $\textrm{Tr}_{E}[|\psi^{\prime}_{SE}\rangle\langle\psi^{\prime}_{SE}|]$ lost the initial coherence in $|\psi_{S}\rangle$.
-In the modern language of the resource theories [14], this an example of the resource interconversion, where one resource, coherence in the pointer basis, gets interconverted into another - entanglement  [15, 16, 17, 18, 19, 20].
-Let us remark here that an abstract coherence theory finds its natural application in open systems with pure dephasing evolution as the latter contains a natural, preferred basis.
-However easy the pure state interconversion (2) looks like, the situation gets much more complicated when considering mixed states, especially if one wants to quantify the amount of the generated entanglement.
+In the modern language of the resource theories [14], this is an example of the resource interconversion, where one resource, coherence in the pointer basis, gets interconverted into another - entanglement  [15, 16, 17, 18, 19, 20].
+Let us remark here that an abstract coherence theory naturally applies to open systems with pure dephasing evolution as the latter contains a natural, preferred basis.
+Although the pure state interconversion (2) appears simple, the situation gets much more complicated when considering mixed states, especially if one wants to quantify the amount of the generated entanglement.
 Assuming that the initial state of the system is pure, Roszak et al.
-[4, 5] recently found necessary and sufficient separability conditions for the system-environment split and have developed own operational measure of the system-environment entanglement [6, 7].
+[4, 5] recently found necessary and sufficient separability conditions for the system-environment split and have developed an operational measure of the system-environment entanglement [6, 7].
 However, it is still unknown how to generalize their methods to mixed initial states of the system.
-Previous studies of the conversion from coherence to entanglement showed that the entanglement generated by incoherent operations between coherent state in the system and incoherent pure state in environment is upper bounded by the initial coherence in the system respect to relative entropy measure [17].
+We introduce some terminology for readers not familiar with quantum coherence.
+Roughly speaking, an incoherent state is simply a state that is diagonal with respect to some reference basis, known as the incoherent basis, and an incoherent operation is a quantum operation that maps incoherent states to incoherent states.
+Detailed definitions will be reviewed later.
+Previous studies of the conversion from coherence to entanglement showed that the entanglement generated by incoherent operations between coherent state in the system and incoherent pure state in environment is upper bounded by the initial coherence in the system with respect to relative entropy measure [17].
 However, for given Hamiltonian, the corresponding pure dephasing evolution is not incoherent operation generally and we also take the environmental state as arbitrary, their finding does not hold in our case.
@@
 In this work, we study entanglement generation for mixed states of the system.
-We derive two lower bounds and one upper bound for the system-environment entanglement as measured by the relative entropy of entanglement $E_{r}$ [21, 22].
-Following the simple intuition (2) and its developments [15, 16, 17, 18, 19, 20], the first lower bound is given in terms of the initial coherence of the system with respect to the natural pointer basis defined by the pure dephasing evolution, diminished by the remaining uncertainty about the system after the best measurement on the environment.
-The second lower bound and the upper bound are quantified with the decoherence in the system, following the intuition that entanglement generation requires not only the initial coherence but the decoherence the system after the evolution which is also suggested by (2).
+We derive a lower bound and an upper bound for the system-environment entanglement as measured by the relative entropy of entanglement $E_{r}$ [21, 22].
+Following the simple intuition (2) and its developments [15, 16, 17, 18, 19, 20], the lower bound is given in terms of the initial coherence of the system with respect to the natural pointer basis defined by the pure dephasing evolution, diminished by the remaining uncertainty about the system after the best measurement on the environment.
+The upper bound is quantified with the strength of decoherence in the system, following the intuition that entanglement generation requires not only the initial coherence but the interaction with the environment followed by decoherence which is also suggested by (2).
 Regarding the first lower bound, for two-dimensional central systems, i.e.
 for qubits, we derive a stronger result in given by a closed formula that links entanglement with the distinguishability of the environment states.
-In particular, this allows to study entanglement dynamics for open qubits, understood as the real-time monitoring of the minimum amount of entanglement produced during the evolution.
+In particular, this allows us to study entanglement dynamics for open qubits, understood as the real-time monitoring of the minimum amount of entanglement produced during the evolution.
 As an illustration, we apply our results to the spin-boson model (see e.g.
@@
 [23].
-2 Preliminaries
-Pure dephasing evolution is defined by the following Hamiltonian:
-$H=H_{S}+H_{E}+\sum_{i=0}^{N-1}|i\rangle\langle i|\otimes V_{i}.$ (3)
-$H_{S}$ is the self Hamiltonian of the system which here we assume to be commuting with the interaction, $H_{S}=\sum_{i}\varepsilon_{i}|i\rangle\langle i|$.
-The case of a more general $H_{S}$ is out of the scope of this work as we concentrate on the pure dephasing models.
-$H_{E}$ and $V_{i}$’s are arbitrary environmental observables.
+II Preliminaries
+We consider here the following Hamiltonian:
+$H_{tot}=H_{S}+H_{E}+\sum_{i=0}^{N-1}|i\rangle\langle i|\otimes V_{i}$ (3)
+where $H_{S}$ is the self-Hamiltonian of the system which we assume to commute with the interaction, i.e.
+$H_{S}=\sum_{i}\varepsilon_{i}|i\rangle\langle i|$.
+The whole resulting Hamiltonian $H_{tot}$ then takes the so-called pure dephasing form: $H_{tot}=\sum_{i}|i\rangle\langle i|\otimes(\varepsilon_{i}+V_{i}+H_{E})\equiv\sum_{i}|i\rangle\langle i|\otimes V_{i}^{\prime}$, with $V_{i}^{\prime}\equiv\varepsilon_{i}+V_{i}+H_{E}$.
+This Hamiltonian preserves the pointer basis, i.e.
+$|i\rangle\langle i|,H_{tot}]=0$, thus affecting only the off-diagonal terms, i.e.
+the coherence, via the dephasing noise, hence the name.
+We make this choice for simplicity, since inclusion of a general $H_{S}$ complicates the dynamics immensely, which forces to use approximate methods such as master equations [2, 3] and ultimately the numerical methods.
+This type of a much more advanced and complicated analysis is out of the scope of the current work and is left for a future research.
+Here, $H_{E}$ and $V_{i}$’s are arbitrary environmental observables.
+Eq.
 (3) is the most general Hamiltonian that generates evolution of the type (2).
 The unitary evolution $U$ is given by:
-$U(t)=\sum_{i}|i\rangle\langle i|\otimes U_{E}^{i}(t),\quad U_{E}^{i}(t)=e^{-(\varepsilon_{i}+H_{E}+V_{i})t},$ (4)
-which is the controlled-unitary or measurement-type operation in the quantum information terminology.
+$U(t)=\sum_{i}|i\rangle\langle i|\otimes U_{E}^{i}(t),\quad U_{E}^{i}(t)=e^{-j(\varepsilon_{i}+H_{E}+V_{i})t},$ (4)
+where $j$ is imaginary unit.
+This unitary is also called the controlled-unitary or measurement-type operation in the quantum information terminology.
 The basis $|i\rangle$ is the pointer basis.
@@
 We will review the definition of incoherent operations below.
-We will assume that the system and the environment are initially product, i.e.
+We will assume that the system and the environment are initially in a product state, i.e.
 $\rho_{SE}(0)=\rho_{S}\otimes\rho_{E}$.
@@
 Applying the unitary operator defined in (4), we obtain the following state $\sigma_{SE}$ after the interaction:
-$\sigma_{SE}=U(\rho_{S}\otimes\rho_{E})U^{\dagger}$
-$=\sum_{i,j}|i\rangle\langle i|\rho_{S}|j\rangle\langle j|\otimes U^{i}_{E}\rho_{E}{U^{j}_{E}}^{\dagger}$ (5)
+$\sigma_{SE}=U(\rho_{S}\otimes\rho_{E})U^{\dagger}=\sum_{i,j}|i\rangle\langle i|\rho_{S}|j\rangle\langle j|\otimes U^{i}_{E}\rho_{E}{U^{j}_{E}}^{\dagger}$ (5)
 In order to evaluate the amount of $SE$ entanglement, we will use the relative entropy of entanglement $E_{r}$ [21, 22], defined as:
-$E_{r}(\rho_{SE})=\min_{\xi_{SE}\in Sep}S(\rho_{SE}||\xi_{SE})$ (6)
+$E_{r}(\rho_{SE})=\min_{\xi_{SE}\in\textrm{Sep}}S(\rho_{SE}||\xi_{SE})$ (6)
 where $S(\rho||\xi)=\textrm{Tr}[\rho(\log\rho-\log\xi)]$ is the quantum relative entropy and the minimization is taken over all separable states $\xi_{SE}=\sum_{i}p_{i}\xi_{S}^{i}\otimes\xi_{E}^{i}$.
+Throughout this paper the base used in the logarithm is 2.
 This is a proper measure of entanglement [24], with a clear operational interpretation as the exponential bound on asymptotic probability of error of confusing $\rho_{AB}$ with the nearest separable state via POVM measurements, see e.g.
@@
 $E_{r}(\rho_{AB})\geq\max\{S(\rho_{A})-S(\rho_{AB}),S(\rho_{B})-S(\rho_{AB})\}$ (7)
-where $S(\rho)$ is the von Neumann entropy: $S(\rho)=-\textrm{Tr}[\rho\log\rho]$.
+where $S(\rho)$ is the von Neumann entropy: $S(\rho)=-\textrm{Tr}[\rho\log\rho]$[27].
 We will also need elements of the quantum coherence theory; see e.g.
@@
 We do not specify the incoherent basis for environment because the coherence in the environment does not matter in our analysis as we will show below.
-In the language of the coherence theory, an incoherent state is simply a diagonal state in the incoherent basis, which is of course nothing else but a decohered state in the open systems/quantum foundations terminology.
-We denote by $\mathcal{I}$ the set of all incoherent states.
-In a resource theory, of which coherence theory is an example, an operation which does not create resource is called a free operation respect to that resource.
+In the language of coherence theory, an incoherent state is simply a diagonal state in the incoherent basis, which is of course nothing else but a decohered state in the open systems/quantum foundations terminology.
+In a resource theory, of which coherence theory is an example, an operation which does not create resource is called a free operation with respect to that resource.
 Various free operations for coherence are proposed for different physical motivations [29].
-The most standard one is incoherent operation whose Kraus operators $\{K_{n}\}$ satisfy $\frac{K_{n}\rho K_{n}^{\dagger}}{\textrm{Tr}[K_{n}\rho K_{n}^{\dagger}]}\in\mathcal{I}$ for all $n$ and any incoherent state $\rho\in\mathcal{I}$.
+The most standard one is incoherent operation whose Kraus operators $\{K_{n}\}$ satisfy that $\frac{K_{n}\rho K_{n}^{\dagger}}{\textrm{Tr}[K_{n}\rho K_{n}^{\dagger}]}$ is incoherent for all $n$ and any incoherent state $\rho$.
 One can now define the relative entropy of coherence $C_{r}$ in analogy to the relative entropy of entanglement (6) [28]:
@@
 Out of the possible measures [30], we will use the state fidelity $B(\rho,\sigma)$, also called generalized overlap, defined as:
-$B(\rho,\sigma)=$ $\min_{\textrm{POVM}:\{M_{i}\}}\sum_{i}\sqrt{\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]}$
+$B(\rho,\sigma)=$ $\min_{\textrm{POVM}:\{M_{m}\}}\sum_{m}\sqrt{\textrm{Tr}[\rho M_{m}]\textrm{Tr}[\sigma M_{m}]}$
 $=$ $\textrm{Tr}\sqrt{\rho^{\frac{1}{2}}\sigma\rho^{\frac{1}{2}}}$ (9)
-where $\textrm{Tr}[\rho M_{i}]\textrm{Tr}[\sigma M_{i}]$ is the probability that one obtains the same value $i$ from the measurement on $\rho$ and $\sigma$.
+where $\textrm{Tr}[\rho M_{m}]\textrm{Tr}[\sigma M_{m}]$ is the probability that one obtains the same value $m$ from the measurement on $\rho$ and $\sigma$.
 Vanishing state fidelity is equivalent to a perfect distinguishability of the two states by POVM’s.
@@
 This measure does not have as clear operational meaning as e.g.
-the quantum Chernoff bound [30], related similarly to the quantum relative entropy to the asymptotic discrimination error of two states, but it is much easier to work with due to the compact form (2) and it also provides both an upper and a lower bound for the quantum Chernoff information.
-Discriminating states from an ensemble =$\mathcal{E}=\{p_{i},\rho_{i}\}$ is a more difficult task and its efficiency can be quantified by the, so called, accessible information $I_{acc}(\mathcal{E})$.
-It is defined as the maximum of the Shannon (classical) mutual information $I(i:m)$ [31] between the ensemble index $i$ and the measurement result $m$, taken over all POVM’s $\{M_{m}\}$ on $\rho=\sum_{i}p_{i}\rho_{i}$.
-The formal expression is given by:
-$I_{acc}=\max_{POVM:\{M_{m}\}}I(i:m),$ (10)
-where $I(i:m)$ is defined using the conditional probabilities of measuring $m$ while the sate was $\rho_{i}$: $p(m|i)=\textrm{Tr}(M_{m}\rho_{i})$.
+the quantum Chernoff bound [31], related similarly to the quantum relative entropy to the asymptotic discrimination error of two states, but it is much easier to work with due to the compact form (II) and it also provides both an upper and a lower bound for the quantum Chernoff information.
+Discriminating states from an ensemble $\mathcal{E}=\{p_{i},\rho_{i}\}_{i}$ is a more difficult task and its efficiency can be quantified by the, so-called, accessible information $I_{acc}(\mathcal{E})$.
+It is defined as the maximum of the Shannon (classical) mutual information $I(\mathcal{I}:\mathcal{M})$ [32] between the variable $\mathcal{I}$ for the ensemble index $i$ and the variable $\mathcal{M}$ for the measurement result $m$, taken over all POVM’s $\{M_{m}\}$ on $\rho=\sum_{i}p_{i}\rho_{i}$.
+Recall that mutual information is defined as I(X:Y)=H(X)H(X|Y)I(X:Y)=H(X)-H(X|Y) where $H(X)$ is the Shannon entropy and $H(X|Y)$ is the conditional entropy for random variable $X$ and $Y$ [32].
+The formal expression of accessible information is given by:
+$I_{acc}=$ $\max_{POVM:\{M_{m}\}}I(\mathcal{I}:\mathcal{M})=\max_{POVM:\{M_{m}\}}H(\mathcal{I})-H(\mathcal{I}|\mathcal{M}),$ (10)
+where the conditional entropy $H(\mathcal{I}|\mathcal{M})$, associated with the POVM $\{M_{m}\}$ is defined as:
+$H(\mathcal{I}|\mathcal{M})=-\sum_{i,m}p_{i}\textrm{Tr}[\rho_{i}M_{m}]\log\frac{p_{i}\textrm{Tr}[\rho_{i}M_{m}]}{\textrm{Tr}[\sum_{j}p_{j}\rho_{j}M_{m}]}.$ (11)
 For two-element ensembles, the two notions of discrimination are connected by the following Lemma, which will be central for our reasoning:
 Lemma 2.
-[32][Lemma 3.6] Let =$\mathcal{E}=\{p,\rho_{0},(1-p),\rho_{1}\}$ be a two-element ensemble of quantum states of arbitrary finite dimensions. Then:
+[33][Lemma 3.6] Let $\mathcal{E}=\{\{(p,\rho_{0}\},\{(1-p),\rho_{1}\}\}$ be a two-element ensemble of quantum states of arbitrary finite dimensions. Then:
 $I_{acc}(\mathcal{E})\geq h(p)-2\sqrt{p(1-p)}B(\rho_{0},\rho_{1})$ (12)
-where $h$ is the binary Shannon entropy: $h(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$.
-3 Bounds on the system-environment entanglement generation
+where $h$ is the binary Shannon entropy: $h(p)=-p\log p-(1-p)\log(1-p)$.
+III Bounds on the system-environment entanglement generation
 In the pure dephasing, open evolution, described in the Introduction, the system’s information is transferred to the environment through the controlled-unitary interaction (4).
@@
 Indeed, as one can easily prove, the state of the environment in the course of the evolution becomes:
-$\sigma_{E}(t)=\textrm{Tr}_{S}[U(t)\rho_{S}\otimes\rho_{E}U(t)^{\dagger}]=\sum_{i}p_{i}\rho_{i}(t),$ (12)
+$\sigma_{E}(t)=\textrm{Tr}_{S}[\sigma_{SE}(t)]=\textrm{Tr}_{S}[U(t)\rho_{S}\otimes\rho_{E}U(t)^{\dagger}]=\sum_{i}p_{i}\rho_{i}(t),$ (13)
 where $p_{i}=\langle i|\rho_{S}|i\rangle$ are the initial pointer probabilities and the conditional states of the environment $\rho_{i}(t)=U_{E}^{i}(t)\rho_{E}{U_{E}^{i}}(t)^{\dagger}$ encode, in a generally imperfect way, the information about the state of the central system.
-One can now try to recover this information by implementing a POVM measurement $\{M_{m}\}$ on the environment in order to extract the value of $i$, encoded by the environment in the course of the evolution.
+One can now try to recover this information by implementing a POVM measurement $\{M_{m}\}$ on the environment in order to extract the value of $\mathcal{I}$, encoded by the environment in the course of the evolution.
 One may view this as an indirect measurement of the system through the environment, common in the macroscopic world; cf.
@@
 [8, 9].
-The uncertainty of this indirect measurement is represented by Shannnon conditional entropy $H(i|m)$ [31].
+The uncertainty of this indirect measurement is represented by Shannon conditional entropy $H(\mathcal{I}|\mathcal{M})$ [32] as $\mathcal{I}$ represents the variable for index on the right hand side of (13) and $\mathcal{M}$ represents the variable for measurement value.
 From above observations, it is natural to think the less this uncertainty is, the stronger the system and the environment are correlated.
@@
 For any POVM measurement $\{M_{m}\}$ on the environment, the relative entropy of entanglement, generated by the pure dephasing evolution $U$, is bounded from below by:
-$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq C_{r}(\rho_{S})-H(i|m).$ (13)
+$E_{r}(\sigma_{SE})\geq C_{r}(\rho_{S})-H(\mathcal{I}|\mathcal{M}).$ (14)
 Proof.
-Applying Lemma 1 to $\sigma_{SE}=U\rho_{S}\otimes\rho_{E}U^{\dagger}$, we have
-[…]
-$S(\sigma_{E})-S(\sigma_{SE})\}.$ (14)
-We focus on finding a lower bound for $S(\sigma_{E})-S(\sigma_{SE})$ because $S(\sigma_{E})$ can be transformed to the conditional entropy or the overlap of environmental states through lemma 2, leading us to a nice result represented in corollary 1.
-We will also consider a lower bound for $S(\sigma_{S})-S(\sigma_{SE})$ later.
+Applying Lemma 1 to $\sigma_{SE}$, we have
+$E_{r}(\sigma_{SE})\geq\max\{S(\sigma_{S})-S(\sigma_{SE}),S(\sigma_{E})-S(\sigma_{SE})\},$ (15)
+where $\sigma_{S}=\textrm{Tr}_{E}[\sigma_{SE}]$.
+We focus on finding a lower bound for the latter, $S(\sigma_{E})-S(\sigma_{SE})$, because $S(\sigma_{E})$ can be transformed to the conditional entropy or the overlap of environmental states through lemma 2, leading us to a nice result represented in corollary 1.
 We first observe that:
@@
 where we used fact that the unitary evolution preserves the von Neumann entropy and it is additive for product states.
-Notice (15) can be negative while the relative entanglement should be always non-negative.
-Next, let us recall the Holevo bound on the $I_{acc}$ associated with (12):
-$I_{acc}\leq S(\sigma_{E})-\sum_{i}\langle i|\rho_{S}|i\rangle S(U_{i}\rho_{E}U_{i}^{\dagger}),$ (16)
+Notice (16) can be negative while the relative entropy of entanglement should be always non-negative.
+Next, to find a lower bound of $S(\sigma_{E})$, let us recall the Holevo bound [34] on the $I_{acc}$ associated with (13):
+$I_{acc}(\sigma_{E})\leq S(\sigma_{E})-\sum_{i}\langle i|\rho_{S}|i\rangle S(U_{i}\rho_{E}U_{i}^{\dagger}),$ (17)
 from which it follows that:
@@
 since $S(U_{i}\rho_{E}U_{i}^{\dagger})=S(\rho_{E})$.
-From the definition of the Shannon mutual information $I(x:y)=H(x)-H(x|y)$, where $H(x)$ and $H(x|y)$ are the classical Shannon and relative entropies respectively, the accessible information $I_{acc}$ can be represented as follows.
-$I_{acc}=$ $\max_{POVM:\{M_{m}\}}I(i:m)$
-$=$ $\max_{POVM:\{M_{m}\}}\left[H(i)-H(i|m)\right].$ (18)
-Index $i$ is a random variable, distributed with the initial pointer probabilities $p_{i}$, as follows from (12), or equivalently with the eigenvalues of the fully decohered state $\Delta(\rho_{S})=\sum_{i}|i\rangle\langle i|\rho_{S}|i\rangle\langle i|$.
-Therefore we may write $H(i)=S(\Delta(\rho_{S}))$.
+From the definition (10), $I_{acc}(\sigma_{E})$ is also bounded by $H(\mathcal{I})-H(\mathcal{I}|\mathcal{M})$.
+$\mathcal{I}$ is distributed with the initial pointer probabilities $p_{i}$, as follows from (13), or equivalently with the eigenvalues of the fully decohered state of system $\Delta(\rho_{S})=\sum_{i}|i\rangle\langle i|\rho_{S}|i\rangle\langle i|$.
+Therefore we may write $H(\mathcal{I})=S(\Delta(\rho_{S}))$.
 Since there is a maximization over all POVM in the definition (10), for an arbitrary POVM $\{M_{m}\}$ we will have:
-$I_{acc}\geq H(i)-H(i|m)=S(\Delta(\rho_{S}))-H(i|m)$ (19)
+$I_{acc}(\sigma_{E})\geq H(\mathcal{I})-H(\mathcal{I}|\mathcal{M})=S(\Delta(\rho_{S}))-H(\mathcal{I}|\mathcal{M})$ (19)
 Now we recover the statement by substituting (18) and (19) into (16):
-$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq$ $S(\sigma_{E})-S(\rho_{S})-S(\rho_{E})$
+$E_{r}(\sigma_{SE})\geq$ $S(\sigma_{E})-S(\rho_{S})-S(\rho_{E})$
 $\geq$ $I_{acc}-S(\rho_{S})$ (20)
-$\geq$ $S(\Delta(\rho_{S}))-H(i|m)-S(\rho_{S})$
-$=$ $C_{r}(\rho_{S})-H(i|m),$
+$\geq$ $S(\Delta(\rho_{S}))-H(\mathcal{I}|\mathcal{M})-S(\rho_{S})$
+$=$ $C_{r}(\rho_{S})-H(\mathcal{I}|\mathcal{M}),$
 where in the final line we used the fact (8). ∎
-The above Lemma is valid for any finete dimensionality of the central system and the environment.
+The above Lemma is valid for any finite dimensionality of the central system and the environment.
 The price to pay for this generality is the presence of the arbitrary POVM in the bound (14).
 This arbitrariness can be removed when the central system is a qubit.
-Then, there ensemble (12) consists of only two states $\rho_{0}(t)=U_{0}(t)\rho_{E}U_{0}(t)^{\dagger}$, $\rho_{1}(t)=U_{1}(t)\rho_{E}U_{1}(t)^{\dagger}$ with probabilities $p=\langle 0|\rho_{S}|0\rangle$ and $1-p=\langle 1|\rho_{S}|1\rangle$.
+Then, their ensemble (13) consists of only two states $\rho_{0}(t)=U_{0}(t)\rho_{E}U_{0}(t)^{\dagger}$, $\rho_{1}(t)=U_{1}(t)\rho_{E}U_{1}(t)^{\dagger}$ with probabilities $p=\langle 0|\rho_{S}|0\rangle$ and $1-p=\langle 1|\rho_{S}|1\rangle$.
 We can then apply the Lemma 2 and obtain:
@@
 If the central system is two-dimensional, then:
-$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})$
-$\geq$ $C_{r}(\rho_{S})-2\sqrt{p(1-p)}B(\rho_{0}(t),\rho_{1}(t)).$ (21)
+$E_{r}(\sigma_{SE})\geq C_{r}(\rho_{S})-2\sqrt{p(1-p)}B(\rho_{0}(t),\rho_{1}(t)).$ (21)
 The proof follows immediately from the application of Lemma 2 in (20).
 Let us look closely at Lemma 3 and give some interpretations.
-The second term $H(i|m)$ on the right hand side is the uncertainty about the system’s state measured by the environment after our best measurement on environment.
-Non zero $H(i|m)$ means that we cannot conclude the system’s value in the pointer basis even if we could use any POVM on the environment.
-Reversely, if $H(i|m)$ becomes zero, the information about the system is transferred fully to environment and entanglement must be there unless system’s coherence is zero, which means the system is fully classical state.
+The second term $H(\mathcal{I}|\mathcal{M})$ on the right hand side is the uncertainty about the system’s state measured by the environment after our best measurement on environment.
+Non zero $H(\mathcal{I}|\mathcal{M})$ means that we cannot conclude the system’s value in the pointer basis even if we could use any POVM on the environment.
+Conversely, if $H(\mathcal{I}|\mathcal{M})$ becomes zero, the information about the system is transferred fully to environment and entanglement must be there unless system’s coherence is zero, which means the system is fully classical state.
 After all, when the system has more coherence in pointer bases, more entanglement is required to know the value of system certainly.
 One can also re-arrange the inequality (14) and obtain the lower bound on the uncertainty about the system’s state in terms of initial coherences and the generated entanglement:
-$\max_{POVM:\{M_{m}\}}H(i|m)\geq C_{r}(\rho_{S})-E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger}).$ (22)
+$\max_{POVM:\{M_{m}\}}H(\mathcal{I}|\mathcal{M})\geq C_{r}(\rho_{S})-E_{r}(\sigma_{SE}).$ (22)
 Similar interpretation can be applied to the qubit version (21).
-If $p(1-p)=0$ one knows the system’s state with certainty.
+If p(1p)=0p(1-p)=0 one knows the system’s state with certainty.
 Notice that $C_{r}(\rho_{S})=0$ in this case.
@@
 This is particularly interesting when the right hand sides of (14),(21) are large, which implies a large amount of created entanglement.
-Finally, we will give an upper bound and another lower bound which we mentioned in the proof of lemma 3 on the generated entanglement.
-From a previous research [17], an upper bound on entanglement generated by incoherent operations with initially coherent system $\rho_{S}$ and an incoherent pure ancilla $|0\rangle\langle 0|_{E}$ is given by the initial coherence in system, for any measure of entanglement $E_{D}$ and coherence $C_{D}$ corresponding to a contractive distance $D$:
+Finally, we will give an upper bound on the generated entanglement.
+From a previous research [17], an upper bound on entanglement generated by incoherent operations with the initially coherent system $\rho_{S}$ and an incoherent pure ancilla $|0\rangle\langle 0|_{E}$ is given by the initial coherence in system, for any measure of entanglement $E_{D}$ and coherence $C_{D}$ corresponding to a contractive divergence $D$:
 $E_{D}(\Lambda[\rho_{S}\otimes|0\rangle\langle 0|_{E}])\leq C_{D}(\rho_{S})$ (23)
 where $\Lambda$ is any incoherent operation on the bipartite system. When we specify $D$ as the relative entropy, which we have been using throughout this work, the above upper bound becomes tight.
-However, since the corresponding unitary evolution $U$ for given Hamiltonian of (3) is not incoherent in general regardless of incoherent basis for environment and we assume the environmental state arbitrary, the above upper bound is not valid in our scenario.
-Also, considering our intuitive picture of system-environment interaction where coherence is converted into entanglement, it is interesting to connect the generated entanglement not only to the initial coherence $C_{r}(\rho_{S})$ as we did in lemma 3 for lower bound, but to the decoherence, the difference of coherence between the initial and final states in system as we demonstrate below.
+However, since the corresponding unitary evolution $U$ for the given Hamiltonian of (3) is not incoherent in general regardless of incoherent basis for environment and we assume the environmental state arbitrary, the above upper bound is not valid in our scenario.
+Also, considering our intuitive picture of system-environment interaction where coherence is converted into entanglement, it is interesting to connect the generated entanglement not only to the initial coherence $C_{r}(\rho_{S})$ as we did in lemma 3 for lower bound, but to the decoherence $C_{r}(\rho_{S})-C_{r}(\sigma_{S})$, the difference of coherence between the initial and final states in system as we demonstrate below.
 Lemma 4.
-Let $D_{S}$ be the decoherence in the system defined as $D_{S}=C_{r}(\rho_{S})-C_{r}(\sigma_{S})$.
-Then the relative entropy of entanglement generated by pure dephasing evolution $U$ is bounded by below and above as:
-$D_{S}-S(\rho_{E})\leq E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\leq D_{S}+H(i).$ (24)
+The relative entropy of entanglement generated by pure dephasing evolution $U$ is bounded from above as:
+$E_{r}(\sigma_{SE})\leq C_{r}(\rho_{S})-C_{r}(\sigma_{S})+H(\mathcal{I}).$ (24)
 Proof.
-Let us prove the left inequality first.
-From lemma 1, $E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq S(\sigma_{S})-S(\sigma_{SE})$.
-The invariancy of entropy over unitary evolution and the fact $S(\Delta(\sigma_{S}))=S(\Delta(\rho_{S}))$ lead us to the followings.
-$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\geq S(\sigma_{S})-S(\sigma_{SE})$
-$=$ $S(\sigma_{S})-S(\rho_{S})-S(\rho_{E})$
-$=$ $S(\sigma_{S})-S(\Delta(\rho_{S}))+S(\Delta(\rho_{S}))-S(\rho_{S})-S(\rho_{E})$
-$=$ $C_{r}(\rho_{S})-C_{r}(\sigma_{S})-S(\rho_{E})$ (25)
-Moving to the right hand side, from the definition of the relative entropy of entanglement: $E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})\leq S(\sigma_{SE}||\sigma_{S}\otimes\sigma_{E})=I(S:E)=S(\sigma_{S})+S(\sigma_{E})-S(\sigma_{SE})=S(\sigma_{S})+S(\sigma_{E})-S(\rho_{S})-S(\rho_{E})$, since by definition $\sigma_{SE}=U\rho_{S}\otimes\rho_{E}U^{\dagger}$ so that $S(\sigma_{SE})=S(\rho_{S}\otimes\rho_{E})=S(\rho_{S})+S(\rho_{E})$.
-We also observe that $S(\sigma_{E})=S(\sum_{i}p_{i}\rho_{i})\leq S(\rho_{E})+H(i)$, which follows from the well known property of the von Neumann entropy: $S(\sum_{i}p_{i}\rho_{i})\leq\sum p_{i}S(\rho_{i})+H(i)$.
+From the definition of the relative entropy of entanglement, $E_{r}(\sigma_{SE})\leq S(\sigma_{SE}||\sigma_{S}\otimes\sigma_{E})$.
+We observe that $\log\sigma_{S}\otimes\sigma_{E}=\log\sigma_{S}\otimes\mathbbm{I}+\mathbbm{I}\otimes\log\sigma_{E}$ and $\textrm{Tr}[\sigma_{SE}\log\sigma_{S}\otimes\mathbbm{I}]=\textrm{Tr}_{S}[\textrm{Tr}_{E}[\sigma_{SE}]\log\sigma_{S}]=\textrm{Tr}_{S}[\sigma_{S}\log\sigma_{S}]=-S(\sigma_{S}).$ Then we have
+$E_{r}(\sigma_{SE})\leq$ $\ S(\sigma_{SE}||\sigma_{S}\otimes\sigma_{E})$
+$=$ $-S(\sigma_{SE})-\textrm{Tr}[\sigma_{SE}\log\sigma_{S}\otimes\sigma_{E}]$
+$=$ $-S(\rho_{S})-S(\rho_{E})-S(\sigma_{S})+S(\sigma_{E})$ (25)
+We also observe that $S(\sigma_{E})\leq S(\rho_{E})+H(\mathcal{I})$, which follows from the well known property of the von Neumann entropy: $S(\sum_{i}p_{i}\rho_{i})\leq\sum p_{i}S(\rho_{i})+H(\mathcal{I})$.
 Substituting this into the above inequality, we prove the statement.
@@
 More precisely, we show that $D_{S}$ is lower bounded by the generated entanglement:
-$E_{r}(U\rho_{S}\otimes\rho_{E}U^{\dagger})-H(i)\leq D_{S}.$ (26)
-4 Case study: The spin-boson model
+$E_{r}(\sigma_{SE})-H(\mathcal{I})\leq C_{r}(\rho_{S})-C_{r}(\sigma_{S}).$ (26)
+IV Case study: The spin-boson model
 In this Section, we illustrate our results on the spin-boson model, which is one of the most popular open systems models; see e.g.
@@
 The dynamics of the model is given by the following Hamiltonian:
-$H$ $=\sum_{\boldsymbol{k}}\omega_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}a_{\boldsymbol{k}}+\sigma_{z}\otimes\sum_{\boldsymbol{k}}\left(g_{\boldsymbol{k}}a^{\dagger}_{\boldsymbol{k}}+g^{*}_{\boldsymbol{k}}a_{\boldsymbol{k}}\right)$ (27)
-where for the purpose of this example, we have suppressed the self-Hamiltonian of the central system, or assumed it to be commuting with the interaction, as otherwise the evolution would not be of the pure dephasing form (4).
-Here, $a^{\dagger}_{\boldsymbol{k}}$, $a_{\boldsymbol{k}}$ are the annihilation and creation operators of the mode $\boldsymbol{k}$ of the environment, $\sigma_{z}$ is the $z$ axis Pauli matrix acting in the space of the central system and $g_{\boldsymbol{k}}$ are the coupling constants.
+$H$ $=\sum_{\bm{k}}\omega_{\bm{k}}a^{\dagger}_{\bm{k}}a_{\bm{k}}+\sigma_{z}\otimes\sum_{\bm{k}}\left(g_{\bm{k}}a^{\dagger}_{\bm{k}}+g^{*}_{\bm{k}}a_{\bm{k}}\right)$ (27)
+where for the purpose of this example, we have suppressed the self-Hamiltonian of the central system, or assumed it commutes with the interaction, as otherwise the evolution would not be of the pure dephasing form (4).
+Here, $a^{\dagger}_{\bm{k}}$, $a_{\bm{k}}$ are the annihilation and creation operators of the mode $\bm{k}$ of the environment, $\sigma_{z}$ is the $z$ axis Pauli matrix acting in the space of the central system and $g_{\bm{k}}$ are the coupling constants.
 The pointer basis here is the eigenbasis of $\sigma_{z}$, which follows immediately from the interaction term in (27).
@@
 $\ln B(t)=$ $\ln B_{vac}(t)+\ln B_{th}(t),$ (31)
-$\frac{1}{2}\ln B_{vac}(t)=$ $-\wp(s-1)\left[1-\frac{\cos[(s-1)\arctan(\Lambda t)]}{(1+\Lambda^{2}t^{2})^{\frac{s-1}{2}}}\right],$ (32)
-$\frac{1}{4}\ln B_{th}(t)=$ $\left(-\frac{T}{2\Lambda}\right)^{s-1}\bigg{\{}\Psi^{(s-2)}\left(1+\frac{T}{2\Lambda}\right)-\Psi^{(s-2)}\left(\frac{1}{2}+\frac{T}{2\Lambda}\right)$
+$\frac{1}{2}\ln B_{vac}(t)=$ $-\Gamma(s-1)\left[1-\frac{\cos[(s-1)\arctan(\Lambda t)]}{(1+\Lambda^{2}t^{2})^{\frac{s-1}{2}}}\right],$ (32)
+$\frac{1}{4}\ln B_{th}(t)=$ $\left(-\frac{T}{2\Lambda}\right)^{s-1}\bigg\{\Psi^{s-2}\left(1+\frac{T}{2\Lambda}\right)-\Psi^{s-2}\left(\frac{1}{2}+\frac{T}{2\Lambda}\right)+\left[\frac{1}{2}\Psi^{s-2}\left(\frac{1}{2}+\frac{T}{2\Lambda}+i\frac{Tt}{2}\right)-\frac{1}{2}\Psi^{s-2}\left(1+\frac{T}{2\Lambda}+i\frac{Tt}{2}\right)+\textrm{c.c.}\right]\bigg\},$ (33)
+where c.c. means complex conjugate of the two terms in the square parenthesis, $\Gamma(s)$ is the Euler gamma function, and $\Psi^{m}(z)$ is the polygamma function defined as:
+$\Psi^{m}(z)=\sum_{k=0}^{\infty}\frac{(-1)^{m+1}m!}{(z+k)^{m+1}}.$ (34)
 […]
-where c.c. means complex conjugate of the last two terms in the parenthesis, $\wp(s)$ is the Euler gamma function, and $\Psi^{m}(z)$ is the polygamma function defined as:
-$\Psi^{m}(z)=\sum_{k=0}^{\infty}\frac{(-1)^{m+1}m!}{(z+k)^{m+1}}.$ (34)
-Figure 1: Lower bound (1) for the spin-boson model as a function of time for a different mixedness parameter $\alpha$ (cf.
-(28)) and different Ohmicity: a) $s=2$ corresponding to Markovian evolution; b) $s=3$ corresponding to non-Markovian evolution.
-The ratio of the thermal to the cutoff energy is set to $T/\Lambda=1$ and the time is measured in the cutoff units.
-The insert shows the peak and thus the non-monotonic behavior of the plots (with different vertical scale for each plot).
-On all plots the parts, where the bound becomes negative were replaced by the trivial bound equal to zero.
-[…]
-[…]
+Figure 2: The same as in Fig. 1 but for a fixed mixedness parameter $\alpha=1/2$ and different ratios $T/\Lambda$. a) Ohmicity $s=2$; b) Ohmicity $s=3$.
+The resulting behavior of the lower bound is presented in Figs.
+1, 2.
+The regions in the plots where the right-hand side of (21) becomes negative, i.e.
+the bound trivializes since by definition $E_{r}\geq 0$ always, were substituted by zero to: i) avoid a suggestion that $E_{r}$ might be negative; ii) emphasize what might be called "a sudden revival of the bound".
+Of course, since this is a lower bound, the fact that in some situations the bound trivializes does not mean there cannot be any entanglement and thus the discontinuity for the initial times need not mean a sudden birth of entanglement.
+But on the other hand, it is intuitively appealing that since we initiate the whole model in a product state, it will take some time for the central system and the environment to get entangled.
 In Fig.
@@
 More interestingly, there is a slight quantitative change of behavior between $s=2$ and $s=3$, corresponding to the transition between the Markovian and non-Markovian evolution respectively.
-While for $s=2$ the behavior is monotonic in time, for $s=3$ there is a broad peak around $\Lambda t=2$, regardless the value of $\alpha$ (depicted on the insert).
+While for $s=2$ the behavior is monotonic in time, for $s=3$ there is a broad peak around $\Lambda t=2$, regardless of the value of $\alpha$ (depicted on the inset).
 This peak is more visible in Fig.
@@
 This in turn leads to a stronger decoherence and explains the well known fact why the system need to be kept cold in order to maintain the coherence.
-5 Conclusions
-We studied generation of entanglement between the system of interest and its environment for a broad class of so called pure dephasing evolutions, also known as measurement-type interactions.
+V Conclusions
+We studied generation of entanglement between the system of interest and its environment for a broad class of so-called pure dephasing evolutions, also known as measurement-type interactions.
 In this part of the work, we extended previous results  [4, 5] obtained for pure initial states of the system, to the case of mixed states.
 We combined methods and results from coherence theory, state discrimination, and open systems evolution to obtain a series of bounds on the amount of the generated entanglement as measured by the relative entropy.
-Firstly, and most importantly, we obtain lower bounds, which certify that at least the given amount of entanglement was created during the evolution out of a given amount of the initial coherence.
+Firstly, and most importantly, we obtained a lower bound, which certifies that at least the given amount of entanglement was created during the evolution out of a given amount of the initial coherence.
 The lower bound character of this result is a price to pay for the generalization to arbitrary, mixed states of the system.
-For qubit as the central system, where we obtained a closed expression, connecting amount of entanglement to the state distinguishability of the environmental states.
+For a qubit as the central system, we obtained a closed expression that connects the amount of entanglement to the state distinguishability of the environmental states.
 This makes an interesting link to certain studies of quantum Darwinism [9].
 For an arbitrary dimension, our bound still contains a certain optimization.
-As an example, we calculated analytically and studied numerically the lower bound for the popular spin-boson system, rediscovering the well-known traces of the non-Markovianity but in the entanglement dynamics.
-Obviously, more systems can be analyzed, especially that for some of them the relevant state distinguishability has already been quantified [37, 38, 39, 33].
-Next, we derived both lower and upper bounds for the entanglement in terms of the difference of coherence in the system before and after the evolution, which is a measure of the decoherence efficiency.
-This proves in an elegant and general way that stronger entanglement leads to more decoherence and vice versa, more decoherence is accompanied with more entanglement generation.
-These bounds also can be seen as an extension of the result obtained in [17].
-Summarizing, our results allow to better characterize and understand the correlations that pure dephasing interaction can produce under general initial conditions.
-Because of the ubiquitous character of the pure dephasing interaction in open systems [2, 3], our results are of a fairly universal character across the field of open systems and decoherence theory.
-Moreover, since quantum measurement theory is largely based on the same type of interaction, we characterize how much entanglement is generated during a measurement process, thus linking to foundations of quantum mechanics.
+As an example, we calculated analytically and numerically the lower bound for the popular spin-boson system, rediscovering well-known traces of the non-Markovianity in the entanglement dynamics.
+Clearly, more systems can be analyzed, especially since for some of them the relevant state distinguishability has already been quantified [39, 40, 41, 35].
+Next, we derived an upper bound for the entanglement in terms of the difference of coherence in the system before and after the evolution, which is a measure of the decoherence efficiency.
+This proves in an elegant and general way that stronger entanglement leads to more decoherence and vice versa, more decoherence is accompanied by more entanglement generation.
+This bound can also be seen as an extension of the result obtained in [17].
+Summarizing, our results allow us to better characterize and understand the correlations that pure dephasing interaction can produce under general initial conditions.
+Because of the ubiquitous character of the pure dephasing interaction in open systems [2, 3], our results are fairly universal across the field of open systems and decoherence theory.
+Moreover, since quantum measurement theory is largely based on the same type of interaction, we characterize how much entanglement is generated during a measurement process, thus linking to the foundations of quantum mechanics.
 On the other hand, a possible practical importance of our studies may lie in the development of non-standard decoherence-mitigation techniques, which rely on the system-environment correlations (see e.g.